### v1

# Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade and with a quantized radiation field studied by the Jaynes-Cummings Model

Abstract
The interaction between atoms and a quantized radiation field is fundamentally important in quantum optics and quantum information science. Due to their unusual properties, Rydberg atoms are promising building blocks for two-qubit gates and atom-light quantum interfaces, exploiting the Rydberg blockade interaction which prevents two atoms at close distance (< 10 $\mu$m) from being simultaneously excited to Rydberg states. Recently, this effect was used to engineer quantum processors based on arrays of interacting Rydberg atoms illuminated by Raman lasers. Motivated by these experiments, we extend the Jaynes-Cummings model to study the interaction between two Rydberg atoms interacting by the Rydberg blockade and a quantized radiation field. We find a rich variety of Rabi oscillations and entanglement between the atoms and the radiation field as a function of initial conditions and interaction time, which may be used to obtain atom-light quantum interfaces as components for future long-distance quantum communication.
1 Introduction
The Jaynes-Cummings model (JCM) is a fundamental model in quantum optics that describes the interaction of a two-level quantum system, such as atoms or other particles, with light under conditions where the quantum nature of photons is significant. Since its introduction in the 1960s [1, 2] and experimental verification in the 1980s [3] (see also [4]), the JCM has been used to describe a wide range of systems involving quantized light-matter interactions, including neutral atoms or ions in an electromagnetic cavity, referred to as cavity quantum electrodynamics (cavity QED) [5, 6], Josephson junctions in superconducting microwave cavities (circuit QED) [7, 8, 9], quantum dots [10, 11, 12, 9], and others; see [13] for a review. Along with this development, the original JCM has been extended in various ways, including intensity-dependent or multiphoton couplings [14, 15, 16, 17, 18], multimode and multi-level atoms [19], many ($N>1$) atoms (Dicke model and Tavis-Cummings model) [20, 21, 22, 23], atoms interacting by dipole-dipole interactions [24, 25, 26], and others [13].
Among the striking properties of quantized light-matter interactions described by the JCM are entanglement between the atoms(s) and the quantized radiation [20, 27, 5, 15, 21, 28, 29, 30], and the collapse and revival of periodic populations of atomic states (Rabi oscillations) for quantum coherent states of the field [31, 13]. These phenomena are entirely due to the quantum nature of the field and do not occur in a semiclassical treatment; indeed, the first experimental verification of field quantization according to the JCM [3, 4] was the observation of the collapse-revival phenomenon predicted in [31].
Entanglement between qubits is a key resource in quantum computation [32]. Therefore, the ability to transform quantum information from stationary qubits encoded in, e.g., trapped atoms, to photons (flying qubits) is a requirement for long-distance quantum communication and the distribution of entanglement over quantum networks [33, 34, 35, 36, 37, 38, 28, 39, 40, 30, 41]. Since the JCM captures the physics of these systems it has been extensively studied in the context of quantum computation and related technologies; see [13, 42] for reviews.
Figure 1: Rydberg blockade effect. Two atoms at close distance (< 10 $\mu$m) in Rydberg states |$\ket{r}$ (dashed cyan circles) interact by van der Waals forces $V_{\text{vdW}}$, resulting in a shift of energy levels as indicated (shown is a case for which $V_{\text{vdW}}$ is positive, which applies to $nS_{1/2}$ Rydberg states; for other Rydberg states $V_{\text{vdW}}$ can be negative [43]). As a result, only one of the two atoms can be excited from the ground state |$\ket{g}$ (filled black dots) to a Rydberg state |$\ket{r}$ (filled cyan dots) by radiation resonant to the frequency $\omega_{0}$ of the transition |$\ket{g}\leftrightarrow\ket{r}$ of an unperturbed atom, but not both atoms simultaneously, if $V_{\text{vdW}}$ is larger than the width of the resonance line. In this case, the doubly excited state |$\ket{r,r}$ is inaccessible to the system.
Figure 2: Schematics of a possible experimental setup for probing two atoms interacting by the Rydberg blockade and with a quantized radiation field in a cavity. The atoms (green balls) are trapped at close distance ($R<10\,\mu$m) by optical tweezers in an electromagnetic cavity and exposed to a radiation field resonant to the atomic transition |$\ket{g}\leftrightarrow\ket{r}$ to excite them to Rydberg states (Figure 1). The laser beams of the optical tweezers are far detuned from the atomic transition.
An almost ideal physical setup for realizing JCM physics are Rydberg atoms interacting with a single mode of a cavity field [44, 4, 45, 28, 46]. Rydberg atoms - atoms (typically alkali) excited to states with large principal quantum numbers $n>20$ (Rydberg states) - interact strongly with one another by van der Waals (vdW) interactions ($V_{\text{vdW}}(R)=C_{6}/R^{6}$ with $C_{6}\sim n^{11}$ where $R$ is the interatomic distance) and with external electric fields (due to their large atomic polarizability $\alpha\sim n^{7}$) [47, 43, 48]. Van der Waals and dipole-dipole energy shifts of pairs of interacting rubidium Rydberg atoms for different quantum numbers $n$, $\ell$, $j$, $m_{j}$ were obtained in [43]. A distinct phenomenon in ensembles of $N\geq 2$ atoms at close distance (< 10 $\mu$m) is the Rydberg blockade effect [49, 50, 51, 52]: Two atoms in a Rydberg state |$\ket{r}$ interact via van der Waals forces over distances of several $\mu$m, whereas the interaction between atoms in the ground state |$\ket{g}$ is negligible; as a result, only one single atom within the blockade radius can be optically excited to a Rydberg state |$\ket{r}$ by radiation resonant to the transition |$\ket{g}\leftrightarrow\ket{r}$ of an unperturbed atom, because the van der Waals interaction shifts the energy levels of all other atoms within the blockade radius out of resonance (illustrated in Figure 1 for two atoms). The Rydberg blockade effect was proposed theoretically as a way to perform fast quantum gates in [49, 50, 51, 52] and experimentally demonstrated for two atoms individually trapped in optical tweezers at a distance of a few $\mu$m in [53, 54, 55]. A CNOT gate based on the Rydberg blockade effect was experimentally demonstrated in [56]. In these experiments, the laser pulses used to excite the atoms to Rydberg states are described as classical electromagnetic waves with precisely controlled duration and power. Conversely, in this paper we consider the case that the radiation is quantized. Figure 2 shows a possible experimental setup for probing two atoms interacting by the Rydberg blockade and with a quantized radiation field in a cavity  [^1] More generally, the quantized radiation interacting with the atoms can be a cavity field as shown in Figure 2 or, e.g., propagating laser light in a quantum coherent state (Section 4)..
In an ensemble of $N>1$ atoms within the blockade radius, a light mode resonant to the transition |$\ket{g}\leftrightarrow\ket{r}$ of an unperturbed atom couples the $N$-body ground state
|$\ket{G}=\ket{g_{1},\ldots,g_{N}}$ (1)
to an entangled $N$-body state
|$\ket{R}=$ (2)
$\frac{1}{\sqrt{N}}\sum\limits_{j=1}^{N}e^{i{\bf k}\cdot{\bf x}_{j}}\ket{g_{1},\ldots,g_{j-1},r_{j},g_{j+1},\ldots,g_{N}}$
in which one Rydberg excitation is coherently shared between all atoms in the ensemble [57, 58, 59]. In (2) |$\ket{g_{j}}$ and |$\ket{r_{j}}$ denote the ground state and a Rydberg state of atom $j$, respectively, ${\bf x}_{j}$ is the position of atom $j$, and $\bf{k}$ is the wave vector of the light mode. Under conditions of a perfect Rydberg blockade any state with two or more atoms excited to |$\ket{r}$ is inaccessible to the system; in this case, the ensemble of $N$ atoms is considered to be an effective two-level Rydberg “superatom” with ground state |$\ket{G}$ and excited state |$\ket{R}$. However, of course, states other than |$\ket{G}$ and |$\ket{R}$ are possible, depending on the initial conditions of the system; for example, the states |$\ket{r,g}$ and |$\ket{g,r}$ can occur in the case of two atoms (Section 3.3). The collective nature of the excitation in |$\ket{R}$ in (2) enhances the coupling of light to a Rydberg superatom by a factor of $\sqrt{N}$; as a result, Rydberg systems with $N\gg 1$ (several hundreds to thousands) atoms have attracted tremendous interest in atomic physics and quantum optics [57, 58, 41, 59]. The Rydberg blockade effect combined with the strong coupling between Rydberg superatoms and light have led to numerous applications of Rydberg atoms in quantum optics and quantum information science, reviewed in [40, 60, 42].
[…]
Motivated by these experiments, in this paper we extend the Jaynes-Cummings model to study two interacting Rydberg atoms under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible (Figure 1). Our motivation is to better understand the interplay between interacting Rydberg atoms and a quantized radiation field in view of the physical implementation of quantum state-transfer between atoms and photons for future quantum information processing [33, 35, 36, 28, 29, 30] and photon-photon quantum gates [66, 67, 68, 41]. In Section 2 we formulate the Hamiltonian of the model and discuss its eigenstates and eigenvalues. We then discuss the dynamics of the system for different initial conditions, including Rabi oscillations and entanglement, for number (Fock) states of the radiation field (Section 3) and a quantum coherent field (Section 4). We conclude in Section 5. Detailed derivations of some results discussed in the main text are left to Appendices A - C.
2 Hamiltonian and eigenstates
In this section, we extend the Jaynes-Cummings model to two Rydberg atoms interacting by the Rydberg blockade. We consider two two-level atoms $a$ and $b$ with ground states |$\ket{g_{a}}$, |$\ket{g_{b}}$ and Rydberg states |$\ket{r_{a}}$, |$\ket{r_{b}}$  [^2] When there is no danger of confusion we omit the labels $a$ and $b$ and denote the ground and Rydberg states for both atoms by |$\ket{g}$ and |$\ket{r}$, respectively., respectively, interacting with a single mode of a quantized radiation field ($f$) with (angular) frequency $\omega_{f}$  [^3] If not stated otherwise, we use the term “frequency” for the angular frequency in units of $\text{rad}\cdot\text{s}^{-1}$.. We assume that the unperturbed transition frequencies $\omega_{0}=\omega_{r}-\omega_{g}$ of the atomic transition |$\ket{g}\leftrightarrow\ket{r}$ of both atoms are equal. The combined system of the two atoms and the quantized field is spanned by product states of the form
[…]
=|$=\ket{\psi_{a},\psi_{b},\phi_{f}}\,.$
We assume that the Rydberg blockade between the atoms is in full effect, so that the doubly excited state, |$\ket{r,r,\phi_{f}}$, is not accessible to the system. Thus, accessible states are spanned by states of the form |$\ket{g,g,n}$, |$\ket{g,r,n}$, and |$\ket{r,g,n}$, with the photon number (Fock) states |$\ket{n}$, $n=0,1,2,\ldots$, of the field. We describe the interaction between the atoms and the quantized field by a Jaynes-Cummings-type Hamiltonian in the rotating wave approximation,  [^4] To simplify notation, in (4) we use the $\otimes$ symbol only to separate operators acting on atoms $a$ and $b$, respectively, but not for the field operators $\hat{a}$, $\hat{a}^{{\dagger}}$.
$\hat{H}$ $=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{0}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (4)
$\,+\hbar\lambda_{a}\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\right)\otimes{\hat{P}}_{g}^{b}$
$\,+{\hat{P}}_{g}^{a}\otimes\hbar\lambda_{b}\left({\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,.$
The first term in (4) is the energy of the field with the photon number operator ${\hat{n}}={\hat{a}}^{\dagger}{\hat{a}}$, where ${\hat{a}}^{\dagger}$, $\hat{a}$ are bosonic creation and annihilation operators and $\hbar=h/\left(2\pi\right)$ is the reduced Planck constant. The second term in (4) is the energy of the unperturbed atoms $a$, $b$ (upper labels) expressed in terms of the Pauli operators ${\hat{\sigma}}_{z}=\mathinner{|{g}\rangle\langle{g}|}-\mathinner{|{r}\rangle\langle{r}|}$, where the zero of energy is chosen as the average of the energies of |$\ket{g}$ and |$\ket{r}$. The interaction terms in the second and third lines of (4) describe processes in which either atom $a$ or atom $b$ is excited from |$\ket{g}$ to |$\ket{r}$ by absorption of one photon, or decays from |$\ket{r}$ to |$\ket{g}$ by emission of one photon, with coupling constants $\lambda_{a}$, $\lambda_{b}$ for atoms $a$, $b$, respectively; ${\hat{\sigma}}_{+}=\mathinner{|{r}\rangle\langle{g}|}$ and ${\hat{\sigma}}_{-}=\mathinner{|{g}\rangle\langle{r}|}$ are raising and lowering Pauli operators for atoms $a$, $b$ (upper labels in (4)). In the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ the Pauli operators are represented by $2\times 2$ matrices  [^5] Note that in the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ (in this order) the raising and lowering Pauli matrices $\sigma_{+}$, $\sigma_{-}$ in (5) are defined the opposite way compared to raising and lowering Pauli matrices for spin-$1/2$ particles.
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&{-1}\end{pmatrix},\,\sigma_{+}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\,\sigma_{-}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\,.$ (5)
Since the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system due to the Rydberg blockade, excitation and decay of one atom can only occur when the other atom is in its ground state |$\ket{g}$; this effect is incorporated in the Hamiltonian (4) by the projection operators ${\hat{P}}_{g}^{a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of atoms $a$ and $b$, respectively.
Following [20] for the case of a stationary field in a cavity, we assume that atom $a$ is located at a peak of the cavity field mode and that the coupling constant $\lambda_{a}$ in (4) is constant and real, i.e.,
$\lambda_{a}\equiv\lambda\in\mathbb{R}\,.$ (6)
The coupling constant $\lambda_{b}$ for atom $b$ then has the form [20]
$\lambda_{b}=\lambda\cos(kR)\,,$ (7)
[…]
$\lambda_{a}=\lambda_{b}=\lambda\in\mathbb{R}\,.$ (8)
Since $\hat{N}={\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b})$ is a constant of motion of the system  [^6] For $\omega_{f}=\omega_{0}$ the quantity $N+1$ corresponds to the total number of energy quanta $\hbar\omega_{f}$ in the system, where $N$ is an eigenvalue of $\hat{N}$., i.e., $[\hat{H},\hat{N}]=0$, it is convenient to write $\hat{H}$ in the form (using (8))
$\hat{H}$ $=\hbar\omega_{f}\hat{N}+\frac{1}{2}\hbar\Delta\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (9)
$\,+\hbar\lambda\Big{[}\Big{(}{\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\Big{)}\otimes{\hat{P}}_{g}^{b}$
$\,+{\hat{P}}_{g}^{a}\otimes\Big{(}{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\Big{)}\Big{]}\,,$
where
$\Delta=\omega_{f}-\omega_{0}$ (10)
is the detuning of the frequency $\omega_{f}$ of the field from the atomic resonance frequency $\omega_{0}$.
Since $\hat{N}$ is a constant of motion, for given photon number $n=0,1,2,\ldots$, the interaction between the atoms and the field in the Hamiltonian (9) couples only the states
|$\ket{\psi_{1}^{(n)}}:=\ket{r,g,n}\,,\,\,\ket{\psi_{2}^{(n)}}:=\ket{g,r,n}\,,\,\,$ (11)
|$\ket{\psi_{3}^{(n)}}:=\ket{g,g,n+1}\,\,.$
This implies that the total Hilbert space $\mathbb{H}$ of the system decouples into subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$, and $\hat{H}$ in (9) can be expressed as a direct sum ${\hat{H}}=\mathop{\oplus}\limits_{n=0}^{\infty}{\hat{H}}^{(n)}$ where the action of each component ${\hat{H}}^{(n)}$ is confined to the subspace $\mathbb{H}_{n}$. In the basis {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ the Hamiltonian ${\hat{H}}^{(n)}$ is represented by a $3\times 3$ matrix
$H^{(n)}=\hbar\begin{pmatrix}\omega_{f}n&0&\lambda\sqrt{n+1}\\[4.0pt] 0&\omega_{f}n&\lambda\sqrt{n+1}\\[4.0pt] \lambda\sqrt{n+1}&\lambda\sqrt{n+1}&\omega_{f}n+\Delta\end{pmatrix}\,\,.$ (12)
For given $n=0,1,2,\ldots$, the energy eigenvalues and associated eigenstates of $H^{(n)}$ are
$E_{\text{asym}\,}^{(n)}\big{/}\hbar=\omega_{f}n\,,$ (13)
$\left\langle\vec{\psi}^{\,(n)}\Big{|}\text{asym},n\right\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ -1\\ 0\end{pmatrix}\,,$ (14)
where $\vec{\psi}^{\,(n)}$ is a vector with components $\psi_{i}^{(n)}$, $i=1,2,3$;
$E_{\pm}^{(n)}\big{/}\hbar=\omega_{f}n+\frac{\Delta}{2}\pm\Omega_{n}(\Delta)\,,$ (15)
$\left\langle\vec{\psi}^{\,(n)}\Big{|}\pm,n\right\rangle=\frac{1}{N_{\pm}}\begin{pmatrix}\pm\lambda\sqrt{n+1}\\[3.0pt] \pm\lambda\sqrt{n+1}\\[2.0pt] \Omega_{n}(\Delta)\pm\frac{\Delta}{2}\end{pmatrix}\,\,.$ (16)
The index “asym” in (14) indicates that this eigenstate is given by the antisymmetric superposition
|$\ket{\text{asym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (17)
$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)\,\,.$
This eigenstate does not involve the basis state |$\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$; as a result, the eigenstate (14) and its eigenvalue (13) are independent of the detuning $\Delta$. In (15) and (16),
$\Omega_{n}(\Delta)=\sqrt{\left(\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$ (18)
is the $\Delta$-dependent Rabi frequency for the two-atom system, which is different from the Rabi frequency $\Omega_{n}(\Delta)$ for one atom [1, 13] by an extra factor of 2 in front of the term $\lambda^{2}(n+1)$ inside the square root. The normalization constants in (16) are $N_{\pm}=\sqrt{\left(\Omega_{n}\pm\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$.
The eigenvectors in (14), (16) form a right-handed, orthonormal frame. This can be made explicit be expressing them in the form
$\left\langle\vec{\psi}^{\,(n)}\Big{|}\text{asym},n\right\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ -1\\ 0\end{pmatrix}\,,\,\,$ (19)
$\left\langle\vec{\psi}^{\,(n)}\Big{|}+,n\right\rangle=\begin{pmatrix}\cos(\varphi_{n})/\sqrt{2}\\[2.0pt] \cos(\varphi_{n})/\sqrt{2}\\[2.0pt] \sin(\varphi_{n})\end{pmatrix}\,,\,\,$ (20)
$\left\langle\vec{\psi}^{\,(n)}\Big{|}-,n\right\rangle=\begin{pmatrix}-\sin(\varphi_{n})/\sqrt{2}\\[2.0pt] -\sin(\varphi_{n})/\sqrt{2}\\[2.0pt] \cos(\varphi_{n})\end{pmatrix}\,,\,$ (21)
corresponding to the column vectors of a $3\times 3$ rotation matrix
$R=\begin{pmatrix}1/\sqrt{2}&\cos(\varphi_{n})/\sqrt{2}&-\sin(\varphi_{n})/\sqrt{2}\\[4.0pt] -1/\sqrt{2}&\cos(\varphi_{n})/\sqrt{2}&-\sin(\varphi_{n})/\sqrt{2}\\[4.0pt] 0&\sin(\varphi_{n})&\cos(\varphi_{n})\end{pmatrix}$ (22)
with an angle
$\varphi_{n}(\Delta)=\tan^{-1}\left(\frac{\Omega_{n}(\Delta)+\Delta/2}{\lambda\sqrt{2(n+1)}}\right)\,\,.$ (23)
Geometrically, $R$ in (22) describes a rotation of the eigenvectors |±$\ket{\pm,n}$ by the angle $\varphi_{n}(\Delta)$ about the fixed axis given by the $\Delta$-independent eigenvector |$\ket{\text{asym},n}$ (Figure 3).
The rotation matrix $R$ in (22) generates the basis transformation {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}\to\left\{\ket{\text{asym},n},\ket{+,n},\ket{-,n}\right\}$ by the relations
|$\ket{\text{asym},n}=\sum\limits_{i=1,2,3}R_{i,\text{asym}}\ket{\psi_{i}^{(n)}}\,,$ (24)
|±$\ket{\pm,n}=\sum\limits_{i=1,2,3}R_{i,\pm}\ket{\psi_{i}^{(n)}}\,.$
The inverse transformation is given by
|$\ket{\psi_{i}^{(n)}}=R_{\text{asym},\,i}^{T}\ket{\text{asym},n}$ (25)
$\quad+R_{+,\,i}^{T}\ket{+,n}+R_{-,\,i}^{T}\ket{-,n}\,,\,\,i=1,2,3\,\,,$
where $R^{T}$ is the transpose of $R$ in (22). In view of the discussion in Section 3, it is useful to display (25) in terms of components:
|$\ket{\psi_{1}^{(n)}}$ $=\frac{1}{\sqrt{2}}\ket{\text{asym},n}+\frac{1}{\sqrt{2}}\cos(\varphi_{n})\ket{+,n}-\frac{1}{\sqrt{2}}\sin(\varphi_{n})\ket{-,n}$ (26)
|$\ket{\psi_{2}^{(n)}}$ $=-\frac{1}{\sqrt{2}}\ket{\text{asym},n}+\frac{1}{\sqrt{2}}\cos(\varphi_{n})\ket{+,n}-\frac{1}{\sqrt{2}}\sin(\varphi_{n})\ket{-,n}$ (27)
|$\ket{\psi_{3}^{(n)}}$ $=\sin(\varphi_{n})\ket{+,n}+\cos(\varphi_{n})\ket{-,n}\,\,,$ (28)
with $\varphi_{n}(\Delta)$ in (23).
Figure 3: Rotation of the eigenvectors |±$\ket{\pm,n}$ (bold green arrows) by the angle $\varphi_{n}(\Delta)$ (magenta) about the fixed axis given by the eigenvector |$\ket{\text{asym},n}$ (bold blue arrow). The rotation is described by the rotation matrix $R$ in (22). The frame is spanned by the states |$\ket{\psi_{i}^{(n)}}$, $i=1,2,3$, in (11). The eigenvectors |±$\ket{\pm,n}$ rotate within the plane perpendicular to |$\ket{\text{asym},n}$ spanned by |$\ket{\text{sym},n}=\tfrac{1}{\sqrt{2}}(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}})$ and |$\ket{\psi_{3}^{(n)}}$ as indicated by the dashed lines. Shown is the configuration of |±$\ket{\pm,n}$ for zero detuning ($\Delta=0$) for which $\varphi_{n}(0)=\pi/4=45^{\circ}$ and $R$ is given by (32).
For zero detuning ($\Delta=0$) the eigenvector in (14) and its eigenvalue in (13) remain unchanged while the remaining expressions simplify to
$\Omega_{n}(0)=\lambda\sqrt{2(n+1)}\,,\,\,\varphi_{n}(0)=\frac{\pi}{4}\,\,,$ (29)
$E_{\pm}^{(n)}\big{/}\hbar=\omega_{f}n\pm\lambda\sqrt{2(n+1)}\,\,,$ (30)
$\left\langle\vec{\psi}^{\,(n)}\Big{|}\pm,n\right\rangle=\begin{pmatrix}\pm 1/2\\[2.0pt] \pm 1/2\\[2.0pt] 1/\sqrt{2}\end{pmatrix}\,\,,$ (31)
$R=\begin{pmatrix}1/\sqrt{2}&1/2&-1/2\\[3.0pt] -1/\sqrt{2}&1/2&-1/2\\[3.0pt] 0&1/\sqrt{2}&1/\sqrt{2}\end{pmatrix}\,\,.$ (32)
Thus, for $\Delta=0$ the states |±$\ket{\pm,n}$ are given by (up to a global phase factor of $-1$ for |$\ket{-,n}$)
|±$\ket{\pm,n}=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{\psi_{3}^{(n)}}\right)$ (33)
$\quad=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{g,g,n+1}\right)\,,$
where (cf. (17), and see Figure 3)
|$\ket{\text{sym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (34)
$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)\,\,.$
The state in (34) corresponds to the state |$\ket{R}$ in (2) for $N=2$ in the absence of the exponential prefactors $e^{i{\bf k}\cdot{\bf x}_{j}}$ (cp. (8)) and for fixed photon number $n$.
3 Number states of the field
In this section, we discuss the dynamics of the system in the subspace $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ (see (11)) for given $n=0,1,2,\ldots$, which is described by the $3\times 3$ matrix $H^{(n)}$ in (12). For a given initial state at $t=0$,
|$\ket{\psi^{(n)}(0)}=\mu_{0}^{(n)}\ket{\psi_{1}^{(n)}}+\nu_{0}^{(n)}\ket{\psi_{2}^{(n)}}+\xi_{0}^{(n)}\ket{\psi_{3}^{(n)}}\,,$ (35)
the dynamics of the system is determined by the time evolution
|$\ket{\psi^{(n)}(t)}=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi^{(n)}(0)}$ (36)
$=\mu^{(n)}(t)\ket{\psi_{1}^{(n)}}+\nu^{(n)}(t)\ket{\psi_{2}^{(n)}}+\xi^{(n)}(t)\ket{\psi_{3}^{(n)}}$
with $\hat{H}$ in (9) and $\mu^{(n)}(t)=\langle\psi_{1}^{(n)}|\psi^{(n)}(t)\rangle$, $\nu^{(n)}(t)=\langle\psi_{2}^{(n)}|\psi^{(n)}(t)\rangle$, $\xi^{(n)}(t)=\langle\psi_{3}^{(n)}|\psi^{(n)}(t)\rangle$. The latter relations can be combined in vector form
$\left\langle\vec{\psi}^{\,(n)}\Big{|}\psi^{(n)}(t)\right\rangle=\begin{pmatrix}\mu^{(n)}(t)\\[2.0pt] \nu^{(n)}(t)\\[2.0pt] \xi^{(n)}(t)\end{pmatrix}\,\,.$ (37)
Depending on the initial state |$\ket{\psi^{(n)}(0)}$ in (35) the time evolution |$\ket{\psi^{(n)}(t)}$ in (36) can be grouped in three general cases A – C discussed in the following subsections.
3.1 Case A: |$\ket{\psi_{A}^{(n)}(0)}=\ket{\text{asym},n}$
with |$\ket{\text{asym},n}=(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}})/\sqrt{2}$ defined in (17). This initial state corresponds to a maximally entangled Bell state of the two atoms and is an eigenstate of $\hat{H}$ in (9) with eigenvalue $E_{\text{asym}}^{(n)}=\hbar\omega_{f}n$ in (13). There is no transition of the field for this state. Thus,
|$\ket{\psi_{A}^{(n)}(t)}=\exp\left(-i\omega_{f}nt\right)\ket{\text{asym},n}$ (38)
which implies for the coefficients in (36)
$\mu_{A}^{(n)}(t)$ $=-\nu_{A}^{(n)}(t)=\frac{1}{\sqrt{2}}\exp\left(-i\omega_{f}nt\right)$ (39)
$\xi_{A}^{(n)}(t)$ $=0\,.$ (40)
The probability $P_{A}^{(n)}(\text{asym},t)$ to find the system in the state |$\ket{\text{asym},n}$ is equal to 1 for all times $t>0$.
For the cases B and C discussed below, we quote the results for the coefficients $\mu^{(n)}(t)$, $\nu^{(n)}(t)$, $\xi^{(n)}(t)$ in (36) and the corresponding probabilities, leaving details to Appendices A and B, respectively. If not stated otherwise, first the result for general detuning $\Delta$ is given, followed by the result for $\Delta=0$.
Figure 4: Amplitude $\sin^{2}(2\varphi_{n})$ of $P_{B}^{(n)}(\text{sym},t)$ in (47), with $\varphi_{n}(\Delta)$ in (23), as a function of the detuning $\Delta$ in units of $\omega_{0}^{-1}$ for $\lambda=\omega_{0}$ and $n=0$ (blue line), 2 (green), and 10 (orange).
3.2 Case B: |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$
In this case, the two atoms are initially in their ground states and are exposed to the field in the photon number state $n+1$ at $t=0$. The coefficients in (36) are given by (see Appendix A.1)
$\mu_{B}^{(n)}(t)$ $=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)$ (41)
$=\frac{1}{\sqrt{2}}(-i)\exp(-i\omega_{f}nt)\sin\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0$ (42)
$\nu_{B}^{(n)}(t)$ $=\mu_{B}^{(n)}(t)$ (43)
$\xi_{B}^{(n)}(t)$ $=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\sin^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\right]$ (44)
$=\exp\Big{(}-i\omega_{f}nt\Big{)}\cos\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0\,\,,$ (45)
with $\varphi_{n}(\Delta)$ in (23). Since $\nu_{B}^{(n)}(t)=\mu_{B}^{(n)}(t)$ the interaction with the field generates the maximally entangled Bell state |$\ket{\text{sym},n}$ for the two atoms in (34). The probability $P_{B}^{(n)}(\text{sym},t)$ to find the system in the state |$\ket{\text{sym},n}$ results as, using
|$\ket{\psi_{B}^{(n)}(t)}=\sigma_{B}^{(n)}(t)\ket{\text{sym},n}+\xi_{B}^{(n)}(t)\ket{\psi_{3}^{(n)}}$ (46)
with $\sigma_{B}^{(n)}(t):=\sqrt{2}\,\mu_{B}^{(n)}(t)$,
$P_{B}^{(n)}(\text{sym},t)$ $=\left|\sigma_{B}^{(n)}(t)\right|^{2}=\sin^{2}(2\varphi_{n})\sin^{2}(\Omega_{n}t)$ (47)
$=\sin^{2}\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0\,\,.$ (48)
The probability to find the system in the complementary state |$\ket{\psi_{3}^{(n)}}$ is given by (see Appendix A.1) $P_{B}^{(n)}\big{(}\ket{\psi_{3}^{(n)}},t\big{)}=1-P_{B}^{(n)}(\text{sym},t)$.
Thus, for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ the system remains in the subspace $\text{span}\{\ket{\text{sym},n},\ket{\psi_{3}^{(n)}}\}$ for all times $t>0$ and oscillates with angular frequency $2\Omega_{n}(\Delta)$ ($=2\lambda\sqrt{2(n+1)}$ for $\Delta=0$) between |$\ket{\psi_{3}^{(n)}}$ and |$\ket{\text{sym},n}$. Since |$\ket{\text{sym},n}$ in (34) corresponds to a maximally entangled Bell state for the two atoms, the interaction of the quantized field with two atoms initially in the ground state generates an entangled state of the atoms that is orthogonal to the entangled state |$\ket{\text{asym},n}$ of case A (Section 3.1).
Due to the amplitude $\sin^{2}(2\varphi_{n})$ in (47), for $\Delta\neq 0$ the value of $P_{B}^{(n)}(\text{sym},t)$ at the maxima is slightly less than 1. For example, for $\lambda=\omega_{0}$, $n=0$, and $\Delta=0.5\,\omega_{0}$, one has $\sin^{2}(2\varphi_{n})\simeq 0.97$. This effect becomes even smaller for increasing $n$ (Figure 4).
In view of Rydberg systems as potential building blocks of atom-light quantum interfaces in quantum networks [36, 39, 41], an important quantity is the entanglement between the two-atom system and the quantized field. Assuming zero detuning for simplicity ($\Delta=0)$, (46) and (42), (45) imply that at times $t$ for which $\lambda\sqrt{2(n+1)}\,t=\pi/4$ the state of the combined system (atoms + field) is given by
|$\ket{\psi_{B}^{(n)}(t)}=\frac{\exp(-i\omega_{f}nt)}{\sqrt{2}}\left(-i\ket{\text{sym},n}+\ket{\psi_{3}^{(n)}}\right)$ (49)
with |$\ket{\text{sym},n}$ in (34) and $\psi_{3}^{(n)}=\ket{g,g,n+1}$ in (11). The state (49) corresponds to a maximally entangled state between the qubit encoded by the two-atom system ($\alpha$) as {|$\left\{\ket{g,g}_{\alpha},\ket{\text{sym}}_{\alpha}\right\}$ and the qubit encoded by the number states of the field ($f$) as {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. If the photons of the field, after interacting with the two-atom system ($\alpha$) resulting in (49), are brought in contact with a second, separate two-atom system ($\beta$) initially prepared in the state |$\ket{g,g}_{\beta}$, then the second two-atom system will evolve to a superposition of |$\ket{g,g}_{\beta}$ and |$\ket{\text{sym}}_{\beta}$ which depends on the photon qubit {|$\left\{\ket{n}_{f},\ket{n+1)}_{f}\right\}$. That is, the second two-atom system ($\beta$) will be entangled with the first ($\alpha$) via the photons of the field. Although this scheme may be challenging to implement experimentally since it requires precise control of the photon number $n$, it illustrates how systems described by the extended Jaynes-Cummings model (4) can in principle be used to entangle stationary atomic systems via photons. In Section 4 we discuss the entanglement between a two-atom system and a quantum single-mode coherent field, which does not require precise control of individual photon number states |$\ket{n}$.
3.3 Case C: |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$
In this case, one of the atoms is initially excited while the other is in the ground state, and the two atoms are exposed to the field in the photon number state |$\ket{n}$ at $t=0$. Since the two atoms are identical, without restriction we consider the case that atom $a$ is excited and atom $b$ is in the ground state at $t=0$. The coefficients in (36) are given by (see Appendix B)
$\begin{matrix}\mu_{C}^{(n)}(t)\\[3.0pt] \nu_{C}^{(n)}(t)\end{matrix}\,\Bigg{\}}$ $=\frac{1}{2}\exp(-i\omega_{f}nt)\Big{[}\pm 1+\exp(-i\tfrac{\Delta}{2}t)\Big{]}\Big{[}\sin^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\cos^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\Big{]}$ (50)
$=\frac{1}{2}\exp(-i\omega_{f}nt)\Big{[}\pm 1+\cos\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\Big{]}\,,\,\,\Delta=0$ (51)
[…]
where the (+)(+)( + ) sign in (50), (51) belongs to $\mu_{C}^{(n)}(t)$ and the ()(-)( - ) sign belongs to $\nu_{C}^{(n)}(t)$. The angle $\varphi_{n}(\Delta)$ is given in (23). For the probabilities of |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ (given by $|\mu_{C}^{(n)}(t)|^{2}$ and $|\nu_{C}^{(n)}(t)|^{2}$, respectively) we only quote explicit formulae for zero detuning ($\Delta=0$) for simplicity; results for $\Delta\neq 0$ are shown in Figure 6 below.
$P_{C}^{(n)}\big{(}\ket{\psi_{1,2}^{(n)}},t\big{)}$ $=\frac{1}{4}\left[1\pm\cos\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\right]^{2}\,,\,\,\Delta=0$ (53)
[…]
In (53) the (+)(+)( + ) sign belongs to |$\ket{\psi_{1}^{(n)}}$ and the ()(-)( - ) sign belongs to |$\ket{\psi_{2}^{(n)}}$.
Figure 5: Time-dependence of the probabilities $P_{C}^{(n)}\big{(}\ket{\psi^{(n)}},t\big{)}$ in (53), (54) for zero detuning ($\Delta=0$) in units of $\Omega_{n}(0)^{-1}=\big{(}\lambda\sqrt{2(n+1)}\,\big{)}^{-1}$, for |$\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ (blue solid line), |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ (green solid line), and |$\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ (orange dotted line).
Figure 5 shows the time-dependence of the probabilities in (53), (54) for $\Delta=0$ in units of $\Omega_{n}(0)^{-1}=\big{(}\lambda\sqrt{2(n+1)}\,\big{)}^{-1}$. The probabilities of the states |$\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ oscillate with angular frequency $\Omega_{n}(0)$ between 0 and 1 and remain completely out of phase at all times. The shape of these probabilities is not purely sinusoidal. Conversely, the probability of |$\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ has a sinusoidal shape and oscillates with angular frequency $2\Omega_{n}(0)$ between 0 and $1/2$.
Figure 6 shows the probabilities corresponding to Figure 5 for finite detuning $\Delta=0.2\,\omega_{0}$. In this case, the time-dependence of the probabilities depends explicitly on $\lambda$ and $\omega_{0}$ as well as on the photon number $n$; in Figure 6, we set $\lambda=\omega_{0}$, express the time-dependence in units of $\omega_{0}^{-1}$, and show results for $n=0$ and $n=10$ (parts A and B). For $\Delta\neq 0$, the relative phase between |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ is no longer constant but oscillates with angular frequency $\Delta$, which for $\Delta\ll\Omega_{n}$ results in an envelope of the waveform that varies slowly compared to the oscillations of the probabilities; the frequency of the latter is an increasing function of $n$. Conversely, the probability of |$\ket{\psi_{3}^{(n)}}$ retains a sinusoidal shape also for nonzero $\Delta$ and oscillates with angular frequency $2\Omega_{n}(\Delta)$ between 0 and $\sin^{2}(2{\varphi_{n}})/2$ (see (54), (47), and Figure 4).
Figure 6: Time-dependence (in units of $\omega_{0}^{-1}$) of the probabilities $P_{C}^{(n)}\big{(}\ket{\psi^{(n)}},t\big{)}$ for |$\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ (blue solid line), |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ (green solid line), and |$\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ (orange dotted line) for $\lambda=\omega_{0}$, detuning $\Delta=0.2\,\omega_{0}$, and (A) $n=0$, (B) $n=10$.
4 Coherent states of the field
In this section we consider the case that the two atoms are initially in their ground states |$\ket{g}$ and the field is in a quantum single-mode coherent state
|$\ket{\alpha}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{m}\,,\,\,t=0\,,$ (55)
where |$\ket{m}$, $m=0,1,2,\ldots$ are photon number states of the field discussed in Section 3, and $\alpha$ is a complex coefficient whose amplitude sets the average number of photons to $\bar{n}=\left|\alpha\right|^{2}$. The quantum coherent state |$\ket{\alpha}$ can describe the radiation field in an electromagnetic cavity (Figure 2) or a propagating laser beam. Thus, the initial state of the system is given by
|$\ket{\psi^{(\alpha)}(0)}$ =|$=\ket{g,g,\alpha}:=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{g,g,m}$ (56)
$=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{g,g,n+1}\right]\,,$
where |$\ket{g,g,n+1}$ are the basis states |$\ket{\psi_{3}^{(n)}}$ of the subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ introduced below (11). Thus, for the Hamiltonian $\hat{H}$ in (9), the initial state |$\ket{\psi^{(\alpha)}(0)}$ in (56) evolves to the state
|$\ket{\psi^{(\alpha)}(t)}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi_{B}^{(n)}(t)}\right]\,,\,\,t>0\,,$ (57)
with |$\ket{\psi_{B}^{(n)}(t)}$ in (46).
We now consider the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\psi_{\text{atoms}}},t\right)$ to find the subsystem of the two atoms at $t>0$ in the state |$\ket{\psi_{\text{atoms}}}$, where the initial state |$\ket{\psi^{(\alpha)}(0)}$ of the total system (atoms + field) is given by (56). Since according to (46) for every $n=0,1,2,\ldots$ the state |$\ket{\psi_{B}^{(n)}(t)}$ in (57) is a linear combination of |$\ket{\text{sym},n}$ and |$\ket{\psi_{3}^{(n)}}$, the state |$\ket{\psi_{\text{atoms}}}$ is contained in the two-dimensional subspace $\mathbb{H}_{\text{atoms}}:=\text{span}\{\ket{\text{sym}},\ket{g,g}\}$, with
|$\ket{\text{sym}}:=\frac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)\,.$ (58)
The probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ to find the two atoms at $t>0$ in the state |$\ket{\text{sym}}$ in (58) is given by the matrix element
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,\text{sym}\rangle\,,$ (59)
where the reduced density operator for the atoms, $\hat{\rho}_{\text{atoms}}(t)$, is given by the partial trace of the density operator
$\hat{\rho}(t)=|\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)|$ (60)
of the total system (atoms + field) in the pure state |$\ket{\psi^{(\alpha)}(t)}$ over the field |$\ket{\phi_{f}}$ (see (3)), i.e.,
$\hat{\rho}_{\text{atoms}}(t)$ $=\text{tr}_{f}\left\{\hat{\rho}(t)\right\}$ (61)
$=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,k\rangle\,,$
where |$\ket{k}$, $k=0,1,2,\ldots$, are the photon number states of the field. Combining (57) - (61) yields
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$
$\quad=\sum\limits_{k=0}^{\infty}\langle\text{sym},k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,\text{sym},k\rangle$
$\quad=\exp(-\bar{n})\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\sigma_{B}^{(m-1)}(t)\right|^{2}$ (62)
$\quad=\exp(-\bar{n})\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\sin^{2}\left(\lambda\sqrt{2m}\,t\right)\,,\,\,\Delta=0\,,$ (63)
with $\sigma_{B}^{(n)}(t):=\sqrt{2}\,\mu_{B}^{(n)}(t)$ in (46). As in (41), (42), first the result for general detuning $\Delta=\omega_{f}-\omega_{0}$ is quoted in (62), followed by the result for $\Delta=0$ in (63). A similar calculation yields for the probability to find the two atoms at $t>0$ in the state |$\ket{g,g}$ the result
$P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=\langle g,g\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle$
$\quad=\exp(-\bar{n})\left(1+\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\xi_{B}^{(m-1)}(t)\right|^{2}\right)$ (64)
$\quad=\exp(-\bar{n})\sum\limits_{m=0}^{\infty}\frac{\bar{n}^{m}}{m!}\cos^{2}\left(\lambda\sqrt{2m}\,t\right)\,,\,\,\Delta=0\,,$ (65)
with $\xi_{B}^{(n)}(t)$ in (44). The normalization condition $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=1$ is fulfilled due to $\left|\sigma_{B}^{n}(t)\right|^{2}+\left|\xi_{B}^{n}(t)\right|^{2}=1$ for $n=0,1,2,\ldots$ (see (46)).
Figure 7 shows $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (blue lines). The oscillations of the probability undergo collapses and revivals on a time scale $\sim\lambda\sqrt{\bar{n}}$, similarly as for a single two-level atom interacting with a quantum single-mode coherent field [31, 13]. However, for the present two-atom system, the oscillations occur between |$\ket{g,g}$ and the maximally entangled Bell state |$\ket{\text{sym}}$ in (58); thus, the interaction of the two-atom system with a quantum single-mode coherent field entangles the two atoms, where the time-dependence of the states and the entanglement can be tuned by the average photon number $\bar{n}$ of the coherent field.
Figure 7: Time-dependence, in units of $\lambda^{-1}$, of the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) (blue lines) and the entanglement $\varepsilon(t)$ in (66) between the two-atom system and the quantum single-mode coherent field (orange lines) for $\omega_{f}=\omega_{0}=\lambda$ and (A) $\bar{n}=10$, (B) $\bar{n}=20$, and (C) $\bar{n}=50$. Part (C) highlights intervals in which $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ first collapses (I) and then remains constant at $1/2$ (II) before the first revival.
In view of the potential use of Rydberg systems for quantum state-transfer between atoms and photons in quantum information processing, it is of interest to study the entanglement of the two-atom system and the quantum coherent field. A measure of this entanglement is the quantity
$\varepsilon(t)=1-\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}\,,$ (66)
which varies between 0 when the atomic and field states are completely disentangled (where $\hat{\rho}_{\text{atoms}}$ describes a pure state) and $1/2$ when the atomic and field states are maximally entangled (where both eigenvalues of $\hat{\rho}_{\text{atoms}}$ are equal to $1/2$). In the basis {|$\left\{\ket{\text{sym}},\ket{g,g}\right\}$ $\hat{\rho}_{\text{atoms}}$ has the form
$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)\end{pmatrix}$ (67)
with
[…]
This yields
$\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}=$ (69)
$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+2\left|\gamma(t)\right|^{2}$
(explicit expressions for $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ are given in Appendix C).
Figure 7 shows the atoms-field entanglement $\varepsilon(t)$ in (66) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (orange lines). As expected, $\varepsilon(0)=0$ for the initial state |$\ket{\psi^{(\alpha)}(0)}$ in (56); subsequently, $\varepsilon(t)$ increases to its maximum value $1/2$ as $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (blue lines in Figure 7) collapses to a constant value $1/2$. Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during the period where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ remains constant at $\simeq 1/2$ after the initial collapse, indicating a nontrivial correlation between the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and the atoms-field entanglement $\varepsilon(t)$.
5 Summary and outlook
We have extended the Jaynes-Cummings model to study two two-level Rydberg atoms, with ground states |$\ket{g}$ and Rydberg states |$\ket{r}$, interacting with a quantum single-mode radiation field under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible to the atoms (Figures 1, 2). We find the eigenstates and eigenvalues of the Hamiltonian (4) for general detuning $\Delta=\omega_{f}-\omega_{0}$ ((13) - (16)) and $\Delta=0$ ((30) - (31)), where $\omega_{f}$ is the frequency of the field and $\omega_{0}$ is the unperturbed transition frequency |$\ket{g}\leftrightarrow\ket{r}$. The eigenstates include the state |$\ket{\text{asym},n}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)$ in (17) for which the atoms are decoupled from the radiation field.
For number (Fock) states of the field, we find explicit expressions for the dynamics of the combined system (atoms + field) for the initial states |$\ket{g,g,n+1}$ (Section 3.2) and |$\ket{r,g,n}$ (Section 3.3), where $n$ is the number of photons. For the initial state |$\ket{g,g,n+1}$ the system remains in the subspace spanned by |$\ket{\text{sym},n}$ and |$\ket{g,g,n+1}$ for all times $t$, where |$\ket{\text{sym},n}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)$ defined in (34). Conversely, for the initial state |$\ket{r,g,n}$ the system oscillates between |$\ket{r,g,n}$ and |$\ket{g,r,n}$ in a non-harmonic way (see Figure 5 for $\Delta=0$ and Figure 6 for $\Delta\neq 0$), undergoing collapses and revivals for $\Delta\neq 0$ similarly as for the Jaynes-Cummings model for one atom [31, 13]. Moreover, for the initial state |$\ket{g,g,n+1}$ we discuss the entanglement between the two-atom system and the quantized field, and describe a scheme how the extended Jaynes-Cummings model (4) can entangle atomic systems via photons in principle (see (49) and text below).
Finally, we find explicit expressions for the dynamics of the combined system (atoms + field) for the case that the field is in a quantum single-mode coherent state, and discuss the entanglement between the two-atom system and the coherent field in terms of the parameter $\varepsilon(t)$ in (66) (Section 4 and Figure 7). Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during an extended period in which the probability of the atomic state |$\ket{\text{sym}}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ remains constant at the value $1/2$ (Figure 7). The non-trivial correlation between the atomic states and the atom-field entanglement $\varepsilon(t)$ is potentially relevant for quantum state-transfer between atoms and photons in quantum information processing.
Our contribution aims to discuss the basic phenomenon of two Rydberg atoms interacting with a quantum single-mode electromagnetic field under perfect Rydberg blockade conditions, allowing for an essentially analytical treatment. The model may be extended in various ways to deal with more realistic experimental conditions. Firstly, any experimental atom-cavity system is an open system coupled to the environment by loss channels and/or propagating light fields, which can be described by a Lindblad master equation for the density operator $\hat{\rho}$ of the system [69], non-Hermitian Hamiltonians [70], and input-output theory [71, 72]. Secondly, a more realistic model will include additional atomic states, e.g., hyperfine ground states of 87Rb to store qubits [63, 65], and multiphoton transitions via intermediate states [19, 16]. Furthermore, recent successful experimental realizations of photon-photon quantum gates use electromagnetically induced transparency to map Rydberg interactions to light [67, 41]; a realistic description of light-matter interactions of this type in the quantum regime will eventually merge quantum optics with a many-body quantum theory of condensed matter physics, where the Jaynes-Cummings model as described here may serve as a reference point [39, 13].
6 Acknowledgements
We thank Volker Quetschke for helpful comments and a critical reading of the manuscript. This work was supported by NSF grants OMA-2231387 and PHY-2012172.
Appendix A Derivations for Section 3.2: Case B
In this appendix we first derive results for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ discussed in Section 3.2 (Appendix A.1), and then for the complementary initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2).
A.1 Initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$
Here we derive the coefficients $\mu_{B}^{(n)}(t)$, $\nu_{B}^{(n)}(t)$, $\xi_{B}^{(n)}(t)$ in (41), (43), (44) and the resulting probabilities. Using (28) one obtains
|$\ket{\psi_{B}^{(n)}(t)}$ $=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi_{3}^{(n)}}$ (70)
$=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\left[\sin(\varphi_{n})\ket{+,n}+\cos(\varphi_{n})\ket{-,n}\right]$ (71)
$=\sin(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)\ket{+,n}+\cos(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\ket{-,n}\,,$ (72)
with $\hat{H}$ in (9) and the eigenvalues $E_{\pm}^{(n)}$ in (15). Combining (72) with (20), (21) one obtains
$\mu_{B}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{B}^{(n)}(t)\right\rangle$ (73)
$=\sin(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)\left\langle\psi_{1}^{(n)}\Big{|}+,n\right\rangle+\cos(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\left\langle\psi_{1}^{(n)}\Big{|}-,n\right\rangle$ (74)
$=\frac{1}{\sqrt{2}}\sin(\varphi_{n})\cos(\varphi_{n})\left[\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)-\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\right]$ (75)
$=\frac{1}{\sqrt{2}}\,\frac{1}{2}\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\exp\left(-i\Omega_{n}t\right)-\exp\left(i\Omega_{n}t\right)\right]$ (76)
$=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,,$ (77)
$\nu_{B}^{(n)}(t)$ $=\left\langle\psi_{2}^{(n)}\Big{|}\psi_{B}^{(n)}(t)\right\rangle=\mu_{B}^{(n)}(t)\,,$ (78)
$\xi_{B}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{B}^{(n)}(t)\right\rangle$ (79)
$=\sin(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)\left\langle\psi_{3}^{(n)}\Big{|}+,n\right\rangle+\cos(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\left\langle\psi_{3}^{(n)}\Big{|}-,n\right\rangle$ (80)
$=\sin^{2}(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)+\cos^{2}(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)$ (81)
$=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\sin^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\cos^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,.$ (82)
Equations (77) and (78) imply for the corresponding probabilities
$P_{B}^{(n)}\left(\ket{\psi_{1}^{(n)}},t\right)=P_{B}^{(n)}\left(\ket{\psi_{2}^{(n)}},t\right)=\left|\mu_{B}^{(n)}(t)\right|^{2}=\frac{1}{2}\sin^{2}(2\varphi_{n})\sin^{2}(\Omega_{n}t)\,.$ (83)
The probability of finding the state |$\ket{\text{sym},n}=\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)/\sqrt{2}$ is given by
$P_{B}^{(n)}(\text{sym},t)=2P_{B}^{(n)}\left(\ket{\psi_{1}^{(n)}},t\right)=\sin^{2}(2\varphi_{n})\sin^{2}(\Omega_{n}t)\,,$ (84)
as quoted in (47). The probability of finding the complementary state |$\ket{\psi_{3}^{(n)}}$ can be calculated similarly, with the result
$P_{B}^{(n)}\left(\ket{\psi_{3}^{(n)}},t\right)$ $=\left|\xi_{B}^{(n)}(t)\right|^{2}$ (85)
$=1-\sin^{2}(2\varphi_{n})\sin^{2}(\Omega_{n}t)$ (86)
$=1-P_{B}^{(n)}(\text{sym},t)\,,$ (87)
as quoted in the text below (48).
A.2 Initial state |$\ket{\psi_{\beta}^{(n)}(0)}=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$
For later use (Appendix B), here we quote results for the coefficients $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ according to (36) and the resulting probabilities for the initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}$, which by (46) and (87) is complementary to the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}$ in Appendix A.1. Similarly as in Appendix A.1 one finds
|$\ket{\psi_{\beta}^{(n)}(t)}$ $=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (88)
$=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\left[\cos(\varphi_{n})\ket{+,n}-\sin(\varphi_{n})\ket{-,n}\right]$ (89)
$=\cos(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)\ket{+,n}-\sin(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\ket{-,n}\,.$ (90)
Combining (90) with (20), (21) one obtains
$\mu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (91)
$=\frac{1}{\sqrt{2}}\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\sin^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,,$ (92)
$\nu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{2}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=\mu_{\beta}^{(n)}(t)\,,$ (93)
$\xi_{\beta}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (94)
$=(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,.$ (95)
The probability $P_{\beta}^{(n)}(\text{sym},t)$ of finding the system in the state |$\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ results as, using |$\ket{\psi_{\beta}^{(n)}(t)}=\sqrt{2}\,\mu_{\beta}^{(n)}(t)\ket{\text{sym},n}+\xi_{\beta}^{(n)}(t)\ket{\psi_{3}^{(n)}}$,
$P_{\beta}^{(n)}(\text{sym},t)=2\left|\mu_{\beta}^{(n)}(t)\right|^{2}$ $=1-\sin^{2}(2\varphi_{n})\sin^{2}(\Omega_{n}t)$ (96)
$=1-P_{B}^{(n)}(\text{sym},t)\,,$ (97)
with $P_{B}^{(n)}(\text{sym},t)$ in (84). Thus, the probabilities of finding the state |$\ket{\text{sym},n}$ at time $t$ for the complementary initial states |$\ket{\psi_{3}^{(n)}}$ and |$\ket{\text{sym},n}$ add up to 1 for all $t$. Furthermore, $P_{\beta}^{(n)}\left(\ket{\psi_{3}^{(n)}},t\right)=\left|\xi_{\beta}^{(n)}(t)\right|^{2}=1-P_{\beta}^{(n)}(\text{sym},t)$, as expected by normalization.
Appendix B Derivations for Section 3.3: Case C
For the initial states |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ considered in Section 3.3 and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ the time evolution |$\ket{\psi^{(n)}(t)}$ can be deduced from case A (Section 3.1) and Appendix A.2, using
|$\ket{\psi_{1}^{(n)}}$ $=\frac{1}{\sqrt{2}}\left(\ket{\psi_{\beta}^{(n)}(0)}+\ket{\psi_{A}^{(n)}(0)}\right)$ (98)
|$\ket{\psi_{2}^{(n)}}$ $=\frac{1}{\sqrt{2}}\left(\ket{\psi_{\beta}^{(n)}(0)}-\ket{\psi_{A}^{(n)}(0)}\right)\,,$ (99)
with |$\ket{\psi_{\beta}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2) and |$\ket{\psi_{A}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (Section 3.1). For the initial state |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}$, (98) implies for the coefficients in (36)
$\mu_{C}^{(n)}(t)$ $=\frac{1}{\sqrt{2}}\left[\mu_{\beta}^{(n)}(t)+\mu_{A}^{(n)}(t)\right]$ (100)
$\nu_{C}^{(n)}(t)$ $=\frac{1}{\sqrt{2}}\left[\nu_{\beta}^{(n)}(t)+\nu_{A}^{(n)}(t)\right]$ (101)
[…]
Using $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ in (91) - (95) and $\mu_{A}^{(n)}(t)$, $\nu_{A}^{(n)}(t)$ in (39) results in (50) - (52) quoted in Section 3.3.
Appendix C Derivation of $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ in (68), (69)
$\gamma(t)$ $=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\langle\text{sym}\,|\,\text{tr}_{f}\left\{\hat{\rho}(t)\right\}|\,g,g\rangle$ (103)
$=\sum\limits_{k=0}^{\infty}\left\langle\text{sym},k\,\Big{|}\psi^{(\alpha)}(t)\right\rangle\left\langle\psi^{(\alpha)}(t)\Big{|}\,g,g,k\right\rangle$ (104)
$=\exp\left(-\left|\alpha\right|^{2}\right)\left[\alpha\,\sigma_{B}^{(0)}(t)\exp(-i\omega_{0}t)+\sum\limits_{k=1}^{\infty}\frac{\alpha^{k+1}}{\sqrt{(k+1)!}}\,\sigma_{B}^{(k)}(t)\,\frac{\alpha^{*k}}{\sqrt{k!}}\,\xi_{B}^{*(k-1)}(t)\right]$ (105)
$=\exp\left(-\left|\alpha\right|^{2}\right)\left[\alpha\,\sigma_{B}^{(0)}(t)\exp(-i\omega_{0}t)+\sum\limits_{k=1}^{\infty}\frac{\left|\alpha\right|^{2k}}{k!}\frac{\alpha}{\sqrt{k+1}}\,\sigma_{B}^{(k)}(t)\,\xi_{B}^{*(k-1)}(t)\right]$ (106)
$=\exp\left(-\left|\alpha\right|^{2}\right)\alpha\,(-i)\exp(-i\omega_{0}t)\sum\limits_{k=0}^{\infty}\frac{\left|\alpha\right|^{2k}}{k!}\frac{1}{\sqrt{k+1}}\sin\left(\lambda\sqrt{2(k+1}\,t\right)\cos\left(\lambda\sqrt{2k}\,t\right)\,,\,\,\Delta=0\,,$ (107)
where (106) holds for general detuning $\Delta$ and (107) for $\Delta=0$. The first term in (105) corresponds to the term $k=0$ of the sum, using
$\langle k=0\,|\,\psi^{(\alpha)}(t)\rangle=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\exp\left(i\omega_{0}t\right)\ket{g,g}+\alpha\,\sigma_{B}^{(0)}(t)\ket{\text{sym}}\right]$ (108)
with |$\ket{\psi^{(\alpha)}(t)}$ from (57). The result (107) is obtained by using the expressions for $\Delta=0$ of the coefficients $\sigma_{B}^{(n)}(t)=\sqrt{2}\mu_{B}^{(n)}(t)$ in (42) and $\xi_{B}^{(n)}(t)$ in (45).
Using (67) and the normalization condition $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=1$ one obtains
$\rho_{\text{atoms}}^{2}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+\left|\gamma(t)\right|^{2}&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+\left|\gamma(t)\right|^{2}\end{pmatrix}$ (109)
which yields (69). For $\left|\gamma(t)\right|^{2}$ we only quote the result for $\Delta=0$ (as used in Figure 7) for simplicity, using (107):
$\left|\gamma(t)\right|^{2}=\exp\left(-2\bar{n}\right)\bar{n}\left[\,\sum\limits_{k=0}^{\infty}\frac{\bar{n}^{k}}{k!}\frac{1}{\sqrt{k+1}}\sin\left(\lambda\sqrt{2(k+1}\,t\right)\cos\left(\lambda\sqrt{2k}\,t\right)\right]^{2}\,,\,\,\Delta=0\,,$ (110)
where $\bar{n}=\left|\alpha\right|^{2}$.

### v2

@@
-Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade and with a quantized radiation field studied by the Jaynes-Cummings Model
+Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade studied by the Jaynes-Cummings Model
 Abstract
@@
 Motivated by these experiments, we extend the Jaynes-Cummings model to study the interaction between two Rydberg atoms interacting by the Rydberg blockade and a quantized radiation field.
-We find a rich variety of Rabi oscillations and entanglement between the atoms and the radiation field as a function of initial conditions and interaction time, which may be used to obtain atom-light quantum interfaces as components for future long-distance quantum communication.
+We consider both number (Fock) states of the field and single-mode quantum coherent states.
+In particular, we discuss different types of entanglements between various components of the total system consisting of the two Rydberg-interacting atoms and coherent states of the field, and show that the behavior is significantly different compared to a system with non-interacting atoms corresponding to the two-atom Tavis-Cummings model.
+Our results are relevant in view of atom-light quantum interfaces as components for future long-distance quantum communication.
+keywords:
+Jaynes-Cummings model , cavity QED , Rydberg atoms , light-matter interactions
+journal: Physics Open
+[…]
+[add]organization=Department of Physics and Astronomy, University of Texas Rio Grande Valley, addressline=1 W University Blvd, city=Brownsville, postcode=TX 78520, country=USA
 1 Introduction
@@
 Along with this development, the original JCM has been extended in various ways, including intensity-dependent or multiphoton couplings [14, 15, 16, 17, 18], multimode and multi-level atoms [19], many ($N>1$) atoms (Dicke model and Tavis-Cummings model) [20, 21, 22, 23], atoms interacting by dipole-dipole interactions [24, 25, 26], and others [13].
-Among the striking properties of quantized light-matter interactions described by the JCM are entanglement between the atoms(s) and the quantized radiation [20, 27, 5, 15, 21, 28, 29, 30], and the collapse and revival of periodic populations of atomic states (Rabi oscillations) for quantum coherent states of the field [31, 13].
+Among the striking properties of quantized light-matter interactions described by the JCM are entanglement between the atoms(s) and the quantized radiation [20, 27, 28, 5, 15, 21, 29, 30, 31], and the collapse and revival of periodic populations of atomic states (Rabi oscillations) for quantum coherent states of the field [32, 28, 13].
 These phenomena are entirely due to the quantum nature of the field and do not occur in a semiclassical treatment; indeed, the first experimental verification of field quantization according to the JCM [3, 4] was the observation of the collapse-revival phenomenon predicted in [32].
+We note that in the Dicke model and Tavis-Cummings model, the atoms do not interact directly with one another but may become entangled via their coupling to a common cavity field [20, 27, 28].
 Entanglement between qubits is a key resource in quantum computation [33].
@@
 to an entangled $N$-body state
-|$\ket{R}=$ (2)
-$\frac{1}{\sqrt{N}}\sum\limits_{j=1}^{N}e^{i{\bf k}\cdot{\bf x}_{j}}\ket{g_{1},\ldots,g_{j-1},r_{j},g_{j+1},\ldots,g_{N}}$
-in which one Rydberg excitation is coherently shared between all atoms in the ensemble [57, 58, 59].
+|$\ket{R}=\frac{1}{\sqrt{N}}\sum\limits_{j=1}^{N}e^{i{\bf k}\cdot{\bf x}_{j}}\ket{g_{1},\ldots,g_{j-1},r_{j},g_{j+1},\ldots,g_{N}}$ (2)
+in which one Rydberg excitation is coherently shared between all atoms in the ensemble [58, 59, 60, 61].
 In (2) |$\ket{g_{j}}$ and |$\ket{r_{j}}$ denote the ground state and a Rydberg state of atom $j$, respectively, ${\bf x}_{j}$ is the position of atom $j$, and $\bf{k}$ is the wave vector of the light mode.
@@
 However, of course, states other than |$\ket{G}$ and |$\ket{R}$ are possible, depending on the initial conditions of the system; for example, the states |$\ket{r,g}$ and |$\ket{g,r}$ can occur in the case of two atoms (Section 3.3).
-The collective nature of the excitation in |$\ket{R}$ in (2) enhances the coupling of light to a Rydberg superatom by a factor of $\sqrt{N}$; as a result, Rydberg systems with $N\gg 1$ (several hundreds to thousands) atoms have attracted tremendous interest in atomic physics and quantum optics [57, 58, 41, 59].
+The collective nature of the excitation in |$\ket{R}$ in (2) enhances the coupling of light to a Rydberg superatom by a factor of $\sqrt{N}$; as a result, Rydberg systems with $N\gg 1$ (several hundreds to thousands) atoms have attracted tremendous interest in atomic physics and quantum optics [58, 59, 60, 42, 61].
 The Rydberg blockade effect combined with the strong coupling between Rydberg superatoms and light have led to numerous applications of Rydberg atoms in quantum optics and quantum information science, reviewed in [41, 62, 43].
@@
 Motivated by these experiments, in this paper we extend the Jaynes-Cummings model to study two interacting Rydberg atoms under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible (Figure 1).
+This direct interaction between the atoms makes the model fundamentally different from the Dicke model and Tavis-Cummings model, in which the atoms do not interact directly with one another [20, 27, 28].
 Our motivation is to better understand the interplay between interacting Rydberg atoms and a quantized radiation field in view of the physical implementation of quantum state-transfer between atoms and photons for future quantum information processing [34, 36, 37, 29, 30, 31] and photon-photon quantum gates [68, 69, 70, 42].
 In Section 2 we formulate the Hamiltonian of the model and discuss its eigenstates and eigenvalues.
-We then discuss the dynamics of the system for different initial conditions, including Rabi oscillations and entanglement, for number (Fock) states of the radiation field (Section 3) and a quantum coherent field (Section 4).
-We conclude in Section 5.
-Detailed derivations of some results discussed in the main text are left to Appendices A - C.
+We then discuss the dynamics of the system for different initial conditions for number (Fock) states of the radiation field (Section 3) and a quantum coherent field (Section 4).
+In Section 5, we discuss different types of entanglements between various components of the system.
+In Section 6, we summarize our results and compare the entanglements for a system with Rydberg-interacting atoms with a system of non-interacting atoms corresponding to the Tavis-Cummings model.
+Detailed derivations of some results discussed in the main text are left to A - D.
 2 Hamiltonian and eigenstates
-In this section, we extend the Jaynes-Cummings model to two Rydberg atoms interacting by the Rydberg blockade.
+In this section, we extend the Jaynes-Cummings model for a single atom  [^2] The Jaynes-Cummings model for a single atom is discussed in many textbooks on quantum optics; see, e.g., [71, 72], and the recent review in [13] for a comprehensive list of references.
+to two Rydberg atoms interacting by the Rydberg blockade.
 We consider two two-level atoms $a$ and $b$ with ground states |$\ket{g_{a}}$, |$\ket{g_{b}}$ and Rydberg states |$\ket{r_{a}}$, |$\ket{r_{b}}$  [^3] When there is no danger of confusion we omit the labels $a$ and $b$ and denote the ground and Rydberg states for both atoms by |$\ket{g}$ and |$\ket{r}$, respectively., respectively, interacting with a single mode of a quantized radiation field ($f$) with (angular) frequency $\omega_{f}$  [^4] If not stated otherwise, we use the term “frequency” for the angular frequency in units of $\text{rad}\cdot\text{s}^{-1}$..
@@
 […]
-=|$=\ket{\psi_{a},\psi_{b},\phi_{f}}\,.$
 We assume that the Rydberg blockade between the atoms is in full effect, so that the doubly excited state, |$\ket{r,r,\phi_{f}}$, is not accessible to the system.
 Thus, accessible states are spanned by states of the form |$\ket{g,g,n}$, |$\ket{g,r,n}$, and |$\ket{r,g,n}$, with the photon number (Fock) states |$\ket{n}$, $n=0,1,2,\ldots$, of the field.
-We describe the interaction between the atoms and the quantized field by a Jaynes-Cummings-type Hamiltonian in the rotating wave approximation,  [^4] To simplify notation, in (4) we use the $\otimes$ symbol only to separate operators acting on atoms $a$ and $b$, respectively, but not for the field operators $\hat{a}$, $\hat{a}^{{\dagger}}$.
-$\hat{H}$ $=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{0}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (4)
-$\,+\hbar\lambda_{a}\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\right)\otimes{\hat{P}}_{g}^{b}$
-$\,+{\hat{P}}_{g}^{a}\otimes\hbar\lambda_{b}\left({\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,.$
+We describe the interaction between the atoms and the quantized field by a Jaynes-Cummings-type Hamiltonian in the rotating wave approximation,  [^5] To simplify notation, in (4) we use the $\otimes$ symbol only to separate operators acting on atoms $a$ and $b$, but not for the field operators $\hat{a}$, $\hat{a}^{{\dagger}}$.
+$\hat{H}=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{0}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda_{a}\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\right)\otimes{\hat{P}}_{g}^{\,b}\,+{\hat{P}}_{g}^{\,a}\otimes\hbar\lambda_{b}\left({\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,.$ (4)
 The first term in (4) is the energy of the field with the photon number operator ${\hat{n}}={\hat{a}}^{\dagger}{\hat{a}}$, where ${\hat{a}}^{\dagger}$, $\hat{a}$ are bosonic creation and annihilation operators and $\hbar=h/\left(2\pi\right)$ is the reduced Planck constant.
 The second term in (4) is the energy of the unperturbed atoms $a$, $b$ (upper labels) expressed in terms of the Pauli operators ${\hat{\sigma}}_{z}=\mathinner{|{g}\rangle\langle{g}|}-\mathinner{|{r}\rangle\langle{r}|}$, where the zero of energy is chosen as the average of the energies of |$\ket{g}$ and |$\ket{r}$.
-The interaction terms in the second and third lines of (4) describe processes in which either atom $a$ or atom $b$ is excited from |$\ket{g}$ to |$\ket{r}$ by absorption of one photon, or decays from |$\ket{r}$ to |$\ket{g}$ by emission of one photon, with coupling constants $\lambda_{a}$, $\lambda_{b}$ for atoms $a$, $b$, respectively; ${\hat{\sigma}}_{+}=\mathinner{|{r}\rangle\langle{g}|}$ and ${\hat{\sigma}}_{-}=\mathinner{|{g}\rangle\langle{r}|}$ are raising and lowering Pauli operators for atoms $a$, $b$ (upper labels in (4)).
+The third and fourth terms in (4) describe processes in which either atom $a$ or atom $b$ is excited from |$\ket{g}$ to |$\ket{r}$ by absorption of one photon, or decays from |$\ket{r}$ to |$\ket{g}$ by emission of one photon, with coupling constants $\lambda_{a}$, $\lambda_{b}$ for atoms $a$, $b$, respectively; ${\hat{\sigma}}_{+}=\mathinner{|{r}\rangle\langle{g}|}$ and ${\hat{\sigma}}_{-}=\mathinner{|{g}\rangle\langle{r}|}$ are raising and lowering Pauli operators for atoms $a$, $b$ (upper labels in (4)).
 In the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ the Pauli operators are represented by $2\times 2$ matrices  [^6] Note that in the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ (in this order) the raising and lowering Pauli matrices $\sigma_{+}$, $\sigma_{-}$ in (5) are defined the opposite way compared to raising and lowering Pauli matrices for spin-$1/2$ particles.
@@
 Since the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system due to the Rydberg blockade, excitation and decay of one atom can only occur when the other atom is in its ground state |$\ket{g}$; this effect is incorporated in the Hamiltonian (4) by the projection operators ${\hat{P}}_{g}^{\,a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{\,b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of atoms $a$ and $b$, respectively.
+The fact that the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system creates a direct interaction between the atoms in the Hamiltonian (4), in addition to the interaction of the atoms with the quantized radiation field.
+This makes our model fundamentally different from previous studies using the Dicke model and Tavis-Cummings model, in which the projection operators ${\hat{P}}_{g}^{\,a}$ and ${\hat{P}}_{g}^{\,b}$ are absent and hence the atoms do not interact directly with one another [20, 27, 28, 29] (see Section 6).
 Following [20] for the case of a stationary field in a cavity, we assume that atom $a$ is located at a peak of the cavity field mode and that the coupling constant $\lambda_{a}$ in (4) is constant and real, i.e.,
@@
 Since $\hat{N}={\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b})$ is a constant of motion of the system  [^7] For $\omega_{f}=\omega_{0}$ the quantity $N+1$ corresponds to the total number of energy quanta $\hbar\omega_{f}$ in the system, where $N$ is an eigenvalue of $\hat{N}$., i.e., $[\hat{H},\hat{N}]=0$, it is convenient to write $\hat{H}$ in the form (using (8))
-$\hat{H}$ $=\hbar\omega_{f}\hat{N}+\frac{1}{2}\hbar\Delta\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (9)
-$\,+\hbar\lambda\Big{[}\Big{(}{\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\Big{)}\otimes{\hat{P}}_{g}^{b}$
-$\,+{\hat{P}}_{g}^{a}\otimes\Big{(}{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\Big{)}\Big{]}\,,$
+$\hat{H}=\hbar\omega_{f}\hat{N}+\frac{1}{2}\hbar\Delta\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda\Big{[}\Big{(}{\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\Big{)}\otimes{\hat{P}}_{g}^{\,b}\,+{\hat{P}}_{g}^{\,a}\otimes\Big{(}{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\Big{)}\Big{]}\,,$ (9)
 where
@@
 Since $\hat{N}$ is a constant of motion, for given photon number $n=0,1,2,\ldots$, the interaction between the atoms and the field in the Hamiltonian (9) couples only the states
-|$\ket{\psi_{1}^{(n)}}:=\ket{r,g,n}\,,\,\,\ket{\psi_{2}^{(n)}}:=\ket{g,r,n}\,,\,\,$ (11)
-|$\ket{\psi_{3}^{(n)}}:=\ket{g,g,n+1}\,\,.$
-This implies that the total Hilbert space $\mathbb{H}$ of the system decouples into subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$, and $\hat{H}$ in (9) can be expressed as a direct sum ${\hat{H}}=\mathop{\oplus}\limits_{n=0}^{\infty}{\hat{H}}^{(n)}$ where the action of each component ${\hat{H}}^{(n)}$ is confined to the subspace $\mathbb{H}_{n}$.
+|$\ket{\psi_{1}^{(n)}}:=\ket{r,g,n}\,,\,\,\ket{\psi_{2}^{(n)}}:=\ket{g,r,n}\,,\,\,\ket{\psi_{3}^{(n)}}:=\ket{g,g,n+1}\,.$ (11)
+Thus, the total Hilbert space $\mathbb{H}$ of the system decouples into 3-dimensional subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$, and $\hat{H}$ in (9) can be expressed as a direct sum ${\hat{H}}=\mathop{\oplus}\limits_{n=0}^{\infty}{\hat{H}}^{(n)}$ where the action of each component ${\hat{H}}^{(n)}$ is confined to the subspace $\mathbb{H}_{n}$  [^8] Without a direct interaction between the atoms as in the Tavis-Cummings model, the doubly excited state |$\ket{r,r,\phi_{f}}$ is accessible to the system.
+As a result, the system decouples into 4-dimensional subspaces (rather than 3-dimensional) that include the state |$\ket{r,r,n-1}$ (for $n\geq 1$) in addition to the 3 states in (11).
+See Section 6, D, and the text in the second paragraph below equation (1) in [28]..
 In the basis {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ the Hamiltonian ${\hat{H}}^{(n)}$ is represented by a $3\times 3$ matrix
@@
 The index “asym” in (14) indicates that this eigenstate is given by the antisymmetric superposition
-|$\ket{\text{asym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (17)
-$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)\,\,.$
+|$\ket{\text{asym},n}:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)\,.$ (17)
 This eigenstate does not involve the basis state |$\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$; as a result, the eigenstate (14) and its eigenvalue (13) are independent of the detuning $\Delta$.
@@
 $\Omega_{n}(\Delta)=\sqrt{\left(\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$ (18)
-is the $\Delta$-dependent Rabi frequency for the two-atom system, which is different from the Rabi frequency $\Omega_{n}(\Delta)$ for one atom [1, 13] by an extra factor of 2 in front of the term $\lambda^{2}(n+1)$ inside the square root.
-The normalization constants in (16) are $N_{\pm}=\sqrt{\left(\Omega_{n}\pm\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$.
+is the $\Delta$-dependent Rabi frequency for the two-atom system. The normalization constants in (16) are
+$N_{\pm}=\sqrt{\left(\Omega_{n}\pm\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$.
 The eigenvectors in (14), (16) form a right-handed, orthonormal frame. This can be made explicit be expressing them in the form
@@
 $\varphi_{n}(\Delta)=\tan^{-1}\left(\frac{\Omega_{n}(\Delta)+\Delta/2}{\lambda\sqrt{2(n+1)}}\right)\,\,.$ (23)
-Geometrically, $R$ in (22) describes a rotation of the eigenvectors |±$\ket{\pm,n}$ by the angle $\varphi_{n}(\Delta)$ about the fixed axis given by the $\Delta$-independent eigenvector |$\ket{\text{asym},n}$ (Figure 3).
+Geometrically, $R$ in (22) describes a rotation of the eigenvectors |±$\ket{\pm,n}$ by the angle $\varphi_{n}(\Delta)$ about the fixed axis given by the $\Delta$-independent eigenvector |$\ket{\text{asym},n}$ (see Figure 3).
 The rotation matrix $R$ in (22) generates the basis transformation {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}\to\left\{\ket{\text{asym},n},\ket{+,n},\ket{-,n}\right\}$ by the relations
 |$\ket{\text{asym},n}=\sum\limits_{i=1,2,3}R_{i,\text{asym}}\ket{\psi_{i}^{(n)}}\,,$ (24)
-|±$\ket{\pm,n}=\sum\limits_{i=1,2,3}R_{i,\pm}\ket{\psi_{i}^{(n)}}\,.$
+|±$\ket{\pm,n}=\sum\limits_{i=1,2,3}R_{i,\pm}\ket{\psi_{i}^{(n)}}\,.$ (25)
 The inverse transformation is given by
-|$\ket{\psi_{i}^{(n)}}=R_{\text{asym},\,i}^{T}\ket{\text{asym},n}$ (25)
-$\quad+R_{+,\,i}^{T}\ket{+,n}+R_{-,\,i}^{T}\ket{-,n}\,,\,\,i=1,2,3\,\,,$
+|$\ket{\psi_{i}^{(n)}}=R_{\text{asym},\,i}^{T}\ket{\text{asym},n}+R_{+,\,i}^{T}\ket{+,n}+R_{-,\,i}^{T}\ket{-,n}\,,\,\,i=1,2,3\,,$ (26)
 where $R^{T}$ is the transpose of $R$ in (22). In view of the discussion in Section 3, it is useful to display (26) in terms of components:
@@
 Thus, for $\Delta=0$ the states |±$\ket{\pm,n}$ are given by (up to a global phase factor of $-1$ for |$\ket{-,n}$)
-|±$\ket{\pm,n}=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{\psi_{3}^{(n)}}\right)$ (33)
-$\quad=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{g,g,n+1}\right)\,,$
+|±$\ket{\pm,n}=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{\psi_{3}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{g,g,n+1}\right)\,,$ (34)
 where (cf. (17), and see Figure 3)
-|$\ket{\text{sym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (34)
-$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)\,\,.$
+|$\ket{\text{sym},n}:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)\,.$ (35)
 The state in (35) corresponds to the state |$\ket{R}$ in (2) for $N=2$ in the absence of the exponential prefactors $e^{i{\bf k}\cdot{\bf x}_{j}}$ (cp. (8)) and for fixed photon number $n$.
@@
 the dynamics of the system is determined by the time evolution
-|$\ket{\psi^{(n)}(t)}=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi^{(n)}(0)}$ (36)
-$=\mu^{(n)}(t)\ket{\psi_{1}^{(n)}}+\nu^{(n)}(t)\ket{\psi_{2}^{(n)}}+\xi^{(n)}(t)\ket{\psi_{3}^{(n)}}$
+|$\ket{\psi^{(n)}(t)}=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\ket{\psi^{(n)}(0)}=\mu^{(n)}(t)\ket{\psi_{1}^{(n)}}+\nu^{(n)}(t)\ket{\psi_{2}^{(n)}}+\xi^{(n)}(t)\ket{\psi_{3}^{(n)}}$ (37)
 with $\hat{H}$ in (9) and $\mu^{(n)}(t)=\langle\psi_{1}^{(n)}|\psi^{(n)}(t)\rangle$, $\nu^{(n)}(t)=\langle\psi_{2}^{(n)}|\psi^{(n)}(t)\rangle$, $\xi^{(n)}(t)=\langle\psi_{3}^{(n)}|\psi^{(n)}(t)\rangle$.
@@
 This initial state corresponds to a maximally entangled Bell state of the two atoms and is an eigenstate of $\hat{H}$ in (9) with eigenvalue $E_{\text{asym}}^{(n)}=\hbar\omega_{f}n$ in (13).
-There is no transition of the field for this state.
+This atomic state is a dark state that does not couple to the field.
 Thus,
@@
 The probability $P_{A}^{(n)}(\text{asym},t)$ to find the system in the state |$\ket{\text{asym},n}$ is equal to 1 for all times $t>0$.
-For the cases B and C discussed below, we quote the results for the coefficients $\mu^{(n)}(t)$, $\nu^{(n)}(t)$, $\xi^{(n)}(t)$ in (36) and the corresponding probabilities, leaving details to Appendices A and B, respectively.
+For the cases B and C discussed below, we quote the results for the coefficients $\mu^{(n)}(t)$, $\nu^{(n)}(t)$, $\xi^{(n)}(t)$ in (37) and the corresponding probabilities, leaving details to A and B, respectively.
 If not stated otherwise, first the result for general detuning $\Delta$ is given, followed by the result for $\Delta=0$.
@@
 3.2 Case B: |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$
-In this case, the two atoms are initially in their ground states and are exposed to the field in the photon number state $n+1$ at $t=0$. The coefficients in (36) are given by (see Appendix A.1)
-$\mu_{B}^{(n)}(t)$ $=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)$ (41)
+In this case, the two atoms are initially in their ground states and are exposed to the field in the photon number state $n+1$ at $t=0$. The coefficients in (37) are given by (see A.1)
+[…]
 $=\frac{1}{\sqrt{2}}(-i)\exp(-i\omega_{f}nt)\sin\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0$ (43)
 $\nu_{B}^{(n)}(t)$ $=\mu_{B}^{(n)}(t)$ (44)
-$\xi_{B}^{(n)}(t)$ $=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\sin^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\right]$ (44)
+$\xi_{B}^{(n)}(t)$ $=\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\cos^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\sin^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\right]$ (45)
 $=\exp\Big{(}-i\omega_{f}nt\Big{)}\cos\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0\,\,,$ (46)
@@
 $=\sin^{2}\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\,,\,\,\Delta=0\,\,.$ (49)
-The probability to find the system in the complementary state |$\ket{\psi_{3}^{(n)}}$ is given by (see Appendix A.1) $P_{B}^{(n)}\big{(}\ket{\psi_{3}^{(n)}},t\big{)}=1-P_{B}^{(n)}(\text{sym},t)$.
+The probability to find the system in the complementary state |$\ket{\psi_{3}^{(n)}}$ is given by (see A.1) $P_{B}^{(n)}\big{(}\ket{\psi_{3}^{(n)}},t\big{)}=1-P_{B}^{(n)}(\text{sym},t)$.
 Thus, for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ the system remains in the subspace $\text{span}\{\ket{\text{sym},n},\ket{\psi_{3}^{(n)}}\}$ for all times $t>0$ and oscillates with angular frequency $2\Omega_{n}(\Delta)$ ($=2\lambda\sqrt{2(n+1)}$ for $\Delta=0$) between |$\ket{\psi_{3}^{(n)}}$ and |$\ket{\text{sym},n}$.
@@
 This effect becomes even smaller for increasing $n$ (Figure 4).
-In view of Rydberg systems as potential building blocks of atom-light quantum interfaces in quantum networks [36, 39, 41], an important quantity is the entanglement between the two-atom system and the quantized field.
+In view of Rydberg systems as potential building blocks of atom-light quantum interfaces, an important quantity is the entanglement between the two-atom system and the quantized field [37, 40, 42].
 Assuming zero detuning for simplicity ($\Delta=0)$, (47) and (43), (46) imply that at times $t$ for which $\lambda\sqrt{2(n+1)}\,t=\pi/4$ the state of the combined system (atoms + field) is given by
@@
 Although this scheme may be challenging to implement experimentally since it requires precise control of the photon number $n$, it illustrates how systems described by the extended Jaynes-Cummings model (4) can in principle be used to entangle stationary atomic systems via photons.
-In Section 4 we discuss the entanglement between a two-atom system and a quantum single-mode coherent field, which does not require precise control of individual photon number states |$\ket{n}$.
+In Section 5 we discuss different types of entanglements between the atoms and a quantum single-mode coherent field, which does not require precise control of individual photon number states |$\ket{n}$.
 3.3 Case C: |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$
 In this case, one of the atoms is initially excited while the other is in the ground state, and the two atoms are exposed to the field in the photon number state |$\ket{n}$ at $t=0$.
-Since the two atoms are identical, without restriction we consider the case that atom $a$ is excited and atom $b$ is in the ground state at $t=0$.
-The coefficients in (36) are given by (see Appendix B)
+Because the Hamiltonian (4) is symmetric under atom exchange, without restriction we assume that atom $a$ is excited and atom $b$ is in the ground state.
+The coefficients in (37) are given by (see B)
 $\begin{matrix}\mu_{C}^{(n)}(t)\\[3.0pt] \nu_{C}^{(n)}(t)\end{matrix}\,\Bigg{\}}$ $=\frac{1}{2}\exp(-i\omega_{f}nt)\Big{[}\pm 1+\exp(-i\tfrac{\Delta}{2}t)\Big{]}\Big{[}\sin^{2}(\varphi_{n})\exp(i\Omega_{n}t)+\cos^{2}(\varphi_{n})\exp(-i\Omega_{n}t)\Big{]}$ (51)
@@
 The angle $\varphi_{n}(\Delta)$ is given in (23).
-For the probabilities of |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ (given by $|\mu_{C}^{(n)}(t)|^{2}$ and $|\nu_{C}^{(n)}(t)|^{2}$, respectively) we only quote explicit formulae for zero detuning ($\Delta=0$) for simplicity; results for $\Delta\neq 0$ are shown in Figure 6 below.
+For the probabilities of |$\ket{\psi_{1}^{(n)}}$ and |$\ket{\psi_{2}^{(n)}}$ (given by $|\hskip 1.0pt\mu_{C}^{(n)}(t)|^{2}$ and $|\nu_{C}^{(n)}(t)|^{2}$, respectively) we only quote explicit formulae for zero detuning ($\Delta=0$) for simplicity; results for $\Delta\neq 0$ are shown in Figure 6 below.
 $P_{C}^{(n)}\big{(}\ket{\psi_{1,2}^{(n)}},t\big{)}$ $=\frac{1}{4}\left[1\pm\cos\Big{(}\lambda\sqrt{2(n+1)}\,t\Big{)}\right]^{2}\,,\,\,\Delta=0$ (54)
@@
 In this section we consider the case that the two atoms are initially in their ground states |$\ket{g}$ and the field is in a quantum single-mode coherent state
-|$\ket{\alpha}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{m}\,,\,\,t=0\,,$ (55)
+|$\ket{\alpha}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{m}\,,\,\,t=0\,,$ (56)
 where |$\ket{m}$, $m=0,1,2,\ldots$ are photon number states of the field discussed in Section 3, and $\alpha$ is a complex coefficient whose amplitude sets the average number of photons to $\bar{n}=\left|\alpha\right|^{2}$.
@@
 Thus, the initial state of the system is given by
-|$\ket{\psi^{(\alpha)}(0)}$ =|$=\ket{g,g,\alpha}:=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{g,g,m}$ (56)
-$=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{g,g,n+1}\right]\,,$
+|$\ket{\psi^{(\alpha)}(0)}$ =|$=\ket{g,g,\alpha}:=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\sum\limits_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}\ket{g,g,m}$ (57)
+$=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{g,g,n+1}\right]\,,$ (58)
 where |$\ket{g,g,n+1}$ are the basis states |$\ket{\psi_{3}^{(n)}}$ of the subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ introduced below (11).
 Thus, for the Hamiltonian $\hat{H}$ in (9), the initial state |$\ket{\psi^{(\alpha)}(0)}$ in (57) evolves to the state
-|$\ket{\psi^{(\alpha)}(t)}=\exp\left(-\tfrac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi_{B}^{(n)}(t)}\right]\,,\,\,t>0\,,$ (57)
+|$\ket{\psi^{(\alpha)}(t)}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi_{B}^{(n)}(t)}\right]\,,\,\,t>0\,,$ (59)
 with |$\ket{\psi_{B}^{(n)}(t)}$ in (47).
@@
 of the total system (atoms + field) in the pure state |$\ket{\psi^{(\alpha)}(t)}$ over the field |$\ket{\phi_{f}}$ (see (3)), i.e.,
-$\hat{\rho}_{\text{atoms}}(t)$ $=\text{tr}_{f}\left\{\hat{\rho}(t)\right\}$ (61)
-$=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,k\rangle\,,$
+$\hat{\rho}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,k\rangle\,,$ (63)
 where |$\ket{k}$, $k=0,1,2,\ldots$, are the photon number states of the field. Combining (59) - (63) yields
-$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$
-$\quad=\sum\limits_{k=0}^{\infty}\langle\text{sym},k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,\text{sym},k\rangle$
+$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)=\sum\limits_{k=0}^{\infty}\langle\text{sym},k\,|\,\psi^{(\alpha)}(t)\rangle\langle\psi^{(\alpha)}(t)\,|\,\text{sym},k\rangle$ (64)
 $\quad=\exp(-\bar{n})\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\sigma_{B}^{(m-1)}(t)\right|^{2}$ (65)
@@
 A similar calculation yields for the probability to find the two atoms at $t>0$ in the state |$\ket{g,g}$ the result
-$P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=\langle g,g\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle$
-$\quad=\exp(-\bar{n})\left(1+\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\xi_{B}^{(m-1)}(t)\right|^{2}\right)$ (64)
+$P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=\langle g,g\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\exp(-\bar{n})\left(1+\sum\limits_{m=1}^{\infty}\frac{\bar{n}^{m}}{m!}\left|\xi_{B}^{(m-1)}(t)\right|^{2}\right)$ (67)
 $\quad=\exp(-\bar{n})\sum\limits_{m=0}^{\infty}\frac{\bar{n}^{m}}{m!}\cos^{2}\left(\lambda\sqrt{2m}\,t\right)\,,\,\,\Delta=0\,,$ (68)
@@
 The normalization condition $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=1$ is fulfilled due to $\left|\sigma_{B}^{n}(t)\right|^{2}+\left|\xi_{B}^{n}(t)\right|^{2}=1$ for $n=0,1,2,\ldots$ (see (47)).
-Figure 7 shows $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (blue lines).
-The oscillations of the probability undergo collapses and revivals on a time scale $\sim\lambda\sqrt{\bar{n}}$, similarly as for a single two-level atom interacting with a quantum single-mode coherent field [31, 13].
+Figure 8 shows $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (61) and $P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)$ in (67) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as functions of time $t$ (orange and gray lines in parts (A), (C), (E), respectively).
+The oscillations of the probabilities undergo collapses and revivals on a time scale $\sim\lambda\sqrt{\bar{n}}$, similarly as for a single two-level atom interacting with a quantum single-mode coherent field [32, 13].
 However, for the present two-atom system, the oscillations occur between |$\ket{g,g}$ and the maximally entangled Bell state |$\ket{\text{sym}}$ in (60); thus, the interaction of the two-atom system with a quantum single-mode coherent field entangles the two atoms, where the time-dependence of the states and the entanglement can be tuned by the average photon number $\bar{n}$ of the coherent field.
-Figure 7: Time-dependence, in units of $\lambda^{-1}$, of the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ in (59) (blue lines) and the entanglement $\varepsilon(t)$ in (66) between the two-atom system and the quantum single-mode coherent field (orange lines) for $\omega_{f}=\omega_{0}=\lambda$ and (A) $\bar{n}=10$, (B) $\bar{n}=20$, and (C) $\bar{n}=50$.
-Part (C) highlights intervals in which $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ first collapses (I) and then remains constant at $1/2$ (II) before the first revival.
-In view of the potential use of Rydberg systems for quantum state-transfer between atoms and photons in quantum information processing, it is of interest to study the entanglement of the two-atom system and the quantum coherent field.
-A measure of this entanglement is the quantity
-$\varepsilon(t)=1-\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}\,,$ (66)
-which varies between 0 when the atomic and field states are completely disentangled (where $\hat{\rho}_{\text{atoms}}$ describes a pure state) and $1/2$ when the atomic and field states are maximally entangled (where both eigenvalues of $\hat{\rho}_{\text{atoms}}$ are equal to $1/2$).
-In the basis {|$\left\{\ket{\text{sym}},\ket{g,g}\right\}$ $\hat{\rho}_{\text{atoms}}$ has the form
-$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)\end{pmatrix}$ (67)
-with
+The phenomenon of entanglement will be discussed in more detail in the following section.
+5 Entanglement
+Figure 7: Entanglements between various components of the total, tripartite system consisting of the two Rydberg-interacting atoms $a$, $b$ (purple) and the quantized radiation field $f$ (red).
+The total system is assumed to be in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59).
+(A) Entanglement $E(ab-f)$ (blue) between the composite two-atom system $ab$ (enclosed by the ellipse) and the field $f$, and entanglement of formation $E(a-b)$ (green) between the two atoms $a$, $b$.
+(B) Entanglement $E(a-bf)$ (magenta) between atom $a$ and the composite (atom $b$)-field system $bf$ (enclosed by the ellipse).
+In view of the potential use of Rydberg systems for quantum state transfer between atoms and photons in quantum information processing [20, 27, 5, 15, 33, 21, 29, 30, 31], it is of interest to study the entanglement between various components of the total system consisting of the two Rydberg-interacting atoms $a$, $b$ and the quantized radiation field (see Figure 7).
+For non-interacting atoms described by the Tavis-Cummings model, corresponding results were obtained in reference [28] (see Section 6).
+To be specific, we assume that the field is in a quantum coherent state |$\ket{\alpha}$ given by (56).
+The time evolution of the entanglement between the atoms and the field depends crucially on the initial condition at $t=0$.
+For example, if the atomic system is initially in the dark state |$\ket{\text{asym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}-\ket{g,r}\right)$ (Section 3.1) the total system (atoms + field) will remain a product state |$\ket{\text{asym}}\otimes\ket{\alpha}$ for all times $t>0$ because the dark state does not couple to the field.
+This implies that the atoms will remain disentangled from the field.
+Conversely, if both atoms are initially in their ground state |$\ket{g}$ (Section 3.2) the atomic system will remain in the two-dimensional subspace perpendicular to |$\ket{\text{asym}}$ and spanned by |$\ket{g,g}$ and |$\ket{\text{sym}}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ for all times $t>0$ (compare Figure 3).
+In this case, the time evolution of the total system (atoms + field) is given by |$\ket{\psi^{(\alpha)}(t)}$ in (59) and discussed in Section 4.
+In what follows, we focus on this case because we expect that the entanglement between the atoms and the field will be strongest when the atomic system is perpendicular to the dark state |$\ket{\text{asym}}$.
+Results for other initial conditions, e.g., one atom excited and the other in the ground state, can be obtained using the results in Section 3.
+5.1 Entanglement between the composite two-atom system $ab$ and the field $f$
 […]
-This yields
-$\text{tr}\{\hat{\rho}_{\text{atoms}}^{2}(t)\}=$ (69)
-$P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+2\left|\gamma(t)\right|^{2}$
-(explicit expressions for $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ are given in Appendix C).
-Figure 7 shows the atoms-field entanglement $\varepsilon(t)$ in (66) for zero detuning ($\Delta=0$) and $\bar{n}=10$, $20$, and $50$ as a function of $t$ (orange lines).
-As expected, $\varepsilon(0)=0$ for the initial state |$\ket{\psi^{(\alpha)}(0)}$ in (56); subsequently, $\varepsilon(t)$ increases to its maximum value $1/2$ as $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (blue lines in Figure 7) collapses to a constant value $1/2$.
-Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during the period where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ remains constant at $\simeq 1/2$ after the initial collapse, indicating a nontrivial correlation between the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and the atoms-field entanglement $\varepsilon(t)$.
-5 Summary and outlook
+If the composite system is in a pure state |$\ket{\psi}$ with density operator $\hat{\rho}=\mathinner{|{\psi}\rangle\langle{\psi}|}$ a quantitative measure for the entanglement $E(A-B)$ between the two subsystems $A$ and $B$ is the von Neumann entropy of either of the two subsystems $A$ and $B$ [33, 73]:
+$E(A-B)=S_{A}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{A}\log_{2}\hat{\rho}_{A}\right\}=S_{B}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{B}\log_{2}\hat{\rho}_{B}\right\}\,,$ (69)
+[…]
+Figure 8: Time-dependence, in units of $\lambda^{-1}$, of the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ in (61) and (67), respectively, and of the entanglements between various components of the total system consisting of the two atoms $a$, $b$ and the quantized radiation field $f$ (Section 5 and Figure 7).
+The total system is initially in the state |$\ket{g,g,\alpha}$ in (57).
+All plots are for $\omega_{f}=\omega_{0}=\lambda$, zero detuning ($\Delta=0$), and for the indicated values of $\bar{n}$.
+(A) Probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (orange lines) and $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ (gray lines) of the atomic states |$\ket{\text{sym}}$ and |$\ket{gg}$, respectively.
+The sum of these probabilities is equal to 1 by normalization.
+(B) Entanglements for $\bar{n}=10$: Entanglement $E(ab-f;t)$ in (70) between the composite two-atom system $ab$ and the field $f$ (blue lines), entanglement $E(a-bf;t)$ in (73) between atom $a$ and the composite (atom $b$)-field system $bf$ (magenta lines), and entanglement of formation $E(a-b;t)$ in (78) between atoms $a$ and $b$ (green lines).
+The colors of the lines in (B) correspond to Figure 7.
+(C), (D) and (E), (F): same as (A), (B) for $\bar{n}=20$ and $\bar{n}=50$, respectively.
+Parts (E) and (F) highlight intervals in which $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ first collapses (I) and then remains constant at $1/2$ (II) before the first revival, showing the close correlation between the atomic probabilities and the entanglements.
+We now apply these definitions to the total system consisting of the two atoms $a$, $b$ and the quantized radiation field $f$.
+We assume that the total system is in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59) with density operator $\hat{\rho}(t)$ in (62) and consider the bipartition $ab\hskip 1.0pt|\hskip 1.0ptf$ of the total system into the composite two-atom system $ab$ (“atoms”) and the field $f$ (see Figure 7).
+Using (69) the entanglement $E(ab-f)$ between the two-atom system $ab$ and the field $f$ is given by
+$E(ab-f)=S_{\text{atoms}}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{\text{atoms}}\log_{2}\hat{\rho}_{\text{atoms}}\right\}\,.$ (70)
+The reduced density operator $\hat{\rho}_{\text{atoms}}=\text{tr}_{f}\left(\hskip 1.0pt\hat{\rho}\right)$ from (63) is represented in the basis {|$\left\{\ket{\text{sym}},\ket{gg}\right\}$ by the matrix
+$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \gamma^{*}(t)&1-P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\hskip 1.0pt\end{pmatrix}\,\,,$ (71)
+where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ from (61) is the probability of finding the two atoms at $t>0$ in the state |$\ket{\text{sym}}$ in (60) and
+[…]
+An explicit expression for $\gamma(t)$ is given in C.
+$E(ab-f)$ in (70) varies between 0 when the atomic and field states are completely disentangled (where $\hat{\rho}_{\text{atoms}}$ describes a pure state) and 1 when the atomic and field states are maximally entangled (where both eigenvalues of $\hat{\rho}_{\text{atoms}}$ are equal to $1/2$).
+Figure 8 shows $E(ab-f)$ in (70) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (blue lines in parts (B), (D), (F), respectively).
+5.2 Entanglement $E(a-bf)$ between atom $a$ and the composite (atom $b$)-field system $bf$
+Following the definitions in Section 5.1 we now consider the bipartition $a\hskip 1.0pt|\hskip 1.0ptbf$ of the total system into atom $a$ and the composite (atom $b$)-field system $bf$ (see Figure 7)  [^9] Switching atoms $a$ and $b$ does not change the results in this subsection because the Hamiltonian (4) is symmetric under atom exchange..
+We again assume that the total system is in the pure state |$\ket{\psi^{(\alpha)}(t)}$ in (59).
+Using (69) the entanglement between the subsystems $a$ and $bf$ is given by
+$E(a-bf)=S_{a}=-\text{tr}\left\{\hskip 1.0pt\hat{\rho}_{a}\log_{2}\hat{\rho}_{a}\right\}\,,$ (73)
+where the reduced density operator $\hat{\rho}_{a}$ of atom $a$ can be obtained as the partial trace of $\hat{\rho}_{\text{atoms}}$ over atom $b$:
+$\hat{\rho}_{a}=\text{tr}_{b,f}\left(\hskip 1.0pt\hat{\rho}\right)=\text{tr}_{b}\left(\hskip 1.0pt\hat{\rho}_{\text{atoms}}\right)\,.$ (74)
+In the basis {|$\left\{\ket{g}_{a},\ket{r}_{a}\right\}$ the density operator $\hat{\rho}_{a}$ is given by the matrix
+$\rho_{a}(t)=\begin{pmatrix}1-\frac{1}{2}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\frac{1}{\sqrt{2}}\gamma^{*}(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\end{pmatrix}\,\,.$ (75)
+Figure 8 shows $E(a-bf)$ in (73) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (magenta lines in parts (B), (D), (F), respectively).
+5.3 Entanglement of formation $E(a-b)$ between the two atoms $a$ and $b$
+Consider again a composite quantum system $Q$ with Hilbert space $\mathbb{H}$ and a bipartition $A|B$ of this system into subsystems $A$ and $B$ as discussed at the beginning of Section 5.1.
+If the composite system is in a mixed state described by a density operator $\hat{\rho}$ rather than in a pure state |$\ket{\psi}$, the entanglement can no longer be obtained as the von Neumann entropy of the subsystems $A$ or $B$.
+Instead, a quantitative measure for the entanglement $E(\hskip 1.0pt\hat{\rho})$ is now given by the entanglement of formation defined as the minimum average entanglement entropy over all pure state decompositions of $\hat{\rho}$, i.e., minimized over all ensembles of states |$\ket{\psi_{i}}$ with probabilities $p_{i}$ such that $\hat{\rho}=\sum_{i}p_{i}\mathinner{|{\psi_{i}}\rangle\langle{\psi_{i}}|}$ [74, 75, 76, 73].
+The entanglement of formation is efficiently computable for two qubits [76], in our case the two two-state atoms $a$ and $b$.
+More general multipartite entanglement measures for pure and mixed states are reviewed in [73].
+[…]
+$E(\hskip 1.0pt\hat{\rho})=h\left(\frac{1}{2}+\frac{1}{2}\sqrt{1-C(\hskip 1.0pt\rho)^{2}}\right)\,\,,$ (76)
+where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$.
+We now apply these definitions to the mixed state of the composite two-atom system $ab$ (“atoms”) represented by the reduced density operator $\hat{\rho}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}$ given in (63).
+This results in the entanglement of formation $E(a-b)$ of the two-atom system $ab$, corresponding to the entanglement of atom $a$ with atom $b$ (see Figure 7).
+In the standard two-qubit basis quoted above the density operator $\hat{\rho}_{\text{atoms}}$ is represented by the matrix
+$\rho_{\text{atoms}}(t)=\begin{pmatrix}P_{gg}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&0\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&0\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&0\\[4.0pt] 0&0&0&0\end{pmatrix}\,\,.$ (77)
+Here, we use the abbreviations $P_{gg}(t)=P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$ and $P_{\text{sym}}(t)=P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ for ease of notation, where $P_{gg}(t)+P_{\text{sym}}(t)=1$ by normalization.
+The rightmost column and the bottom row of the matrix (77) are filled with zeros because the doubly excited state |$\ket{rr}$ is not available to the system due to the Rydberg blockade between the atoms.
+We find that the matrix $\rho_{\text{atoms}}\cdot\tilde{\rho}_{\text{atoms}}$ has only 1 nonzero eigenvalue given by $P_{\text{sym}}^{\,2}$ resulting in $C(\hskip 1.0pt\rho_{\text{atoms}})=P_{\text{sym}}=P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}}\right)$.
+That is, the concurrence is precisely equal to the probability of finding the two atoms in the state |$\ket{\text{sym}}$ in (60).
+Using (76) we obtain the entanglement of formation as a function of time $t$:
+$E(a-b;t)=h\left(\frac{1}{2}+\frac{1}{2}\sqrt{1-\left[P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)\right]^{2}}\right)\,\,.$ (78)
+Clearly, $E(a-b;t)$ is entirely determined by the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ (Figure 9).
+The limiting behaviors are $E(a-b)\sim P_{\text{sym}}^{\,2}$ for $P_{\text{sym}}\to 0$ (with a logarithmic correction) and $E(a-b)\to 1$ linearly for $P_{\text{sym}}\to 1$.
+Figure 9: Entanglement of formation $E(a-b)$ between the two atoms $a$ and $b$ in (78) as a function of the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}}\right)$ of finding the two atoms in the state |$\ket{\text{sym}}$ in (60).
+Figure 8 shows $E(a-b;t)$ in (78) as a function of time $t$ for $\bar{n}=10$, $\bar{n}=20$, and $\bar{n}=50$ (green lines in parts (B), (D), (F), respectively).
+Figure 8 shows that the three forms of entanglement $E(ab-f,t)$, $E(a-bf,t)$, $E(a-b;t)$ (Figure 7) and the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ are strongly correlated.
+The direct correlation between $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ and $E(a-b;t)$ is evident by (78) and Figure 9.
+As expected, all three forms of entanglement vanish for $t=0$ because the total system is initially prepared in the product state |$\ket{g,g,\alpha}$ in (57).
+Subsequently, $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ collapses to a constant value $1/2$ (orange lines in parts (A), (C), (E)).
+Interestingly, $E(ab-f,t)$ and $E(a-bf,t)$ (blue and magenta lines in parts (B), (D), (F)) do not remain constant but decrease to relatively low values during the period where $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ remains constant at $\simeq 1/2$ after the initial collapse.
+6 Summary and conclusions
 We have extended the Jaynes-Cummings model to study two two-level Rydberg atoms, with ground states |$\ket{g}$ and Rydberg states |$\ket{r}$, interacting with a quantum single-mode radiation field under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible to the atoms (Figures 1, 2).
@@
 Moreover, for the initial state |$\ket{g,g,n+1}$ we discuss the entanglement between the two-atom system and the quantized field, and describe a scheme how the extended Jaynes-Cummings model (4) can entangle atomic systems via photons in principle (see (50) and text below).
-Finally, we find explicit expressions for the dynamics of the combined system (atoms + field) for the case that the field is in a quantum single-mode coherent state, and discuss the entanglement between the two-atom system and the coherent field in terms of the parameter $\varepsilon(t)$ in (66) (Section 4 and Figure 7).
-Interestingly, $\varepsilon(t)$ is not constant but decreases to a value close to 0 during an extended period in which the probability of the atomic state |$\ket{\text{sym}}=\tfrac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ remains constant at the value $1/2$ (Figure 7).
-The non-trivial correlation between the atomic states and the atom-field entanglement $\varepsilon(t)$ is potentially relevant for quantum state-transfer between atoms and photons in quantum information processing.
+We find explicit expressions for the dynamics of the combined system (atoms + field) for the case that the field is in a quantum single-mode coherent state (Section 4 and Figure 8).
+Using these results, we obtain the entanglement between various components of the total system consisting of the two Rydberg-interacting atoms $a$, $b$ and the quantized radiation field (Figures 7 and 8).
+We find distinct correlations between these entanglements and the probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ of finding the atoms in the symmetric state |$\ket{\text{sym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$ in (60), which are absent for a system of non-interacting atoms studied by the Tavis-Cummings model (see below).
+The correlations between the atomic states and the atom-field entanglement is potentially relevant for quantum state-transfer between atoms and photons in quantum information processing.
+It is interesting to compare the three types of entanglement for the system with Rydberg-interacting atoms shown in parts (B), (D), and (F) of Figure 8 with the corresponding entanglements for a system with non-interacting atoms described by the two-atom Tavis-Cummings model (TC).
+For zero detuning and identical coupling constant $\lambda$ between each of the atoms $a$ and $b$ and the field, the TC model is governed by the Hamiltonian
+$\hat{H}_{TC}=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{f}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}+{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,,$ (79)
+where $\omega_{f}$ is again the angular frequency of the quantized radiation field, which for zero detuning is equal to the atomic transition frequency $\omega_{0}$.
+The TC Hamiltonian (79) is obtained from the Hamiltonian (4) by replacing the projection operators ${\hat{P}}_{g}^{\,a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{\,b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of the atoms with the identity operator $\hat{I}$, so that the doubly excited state |$\ket{r,r,\phi_{f}}$ is accessible in the TC model.
+Entanglements between the atoms and the field in the TC model were studied in reference [28] in terms of tangles, which is a different (albeit closely related, see D) entanglement measure than the von Neumann entropy and entanglement of formation discussed in Section 5 and shown in our Figure 8.
+To facilitate a one-to-one comparison of the behavior of the system with Rydberg-interacting atoms (this work) and independent atoms (TC model), Figure 10 shows the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$, $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$, and $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ of the respective atomic states |$\ket{\text{sym}}$, |$\ket{gg}$, and |$\ket{rr}$ (see (126) in D), as well as the three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ defined in (70), (73), and (78), respectively, obtained for the TC model in (79) (see D for details).
+Figure 10 should be compared with Figure 8.
+The parameters, labels, and the meaning of the colored lines in Figure 10 are the same as in Figure 8; the only exception are the probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ shown as purple lines in parts (A), (C), (E) in Figure 10, which are absent in Figure 8 because the doubly excited state |$\ket{rr}$ is not accessible for Rydberg-interacting atoms.
+Figures 8 and 10 show that the time dependence of the three types of entanglement in parts (B), (D), and (F) is notably different for Rydberg-interacting atoms compared to the TC model.
+In particular, for the TC model (Figure 10), the entanglements are less correlated with the atomic probability $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$ than for the system with Rydberg-interacting atoms (Figure 8).
+Moreover, the atom-atom entanglement $E(a-b;t)$ is significantly lower for the TC model than for the system with Rydberg-interacting atoms.
+Thus, the Rydberg blockade interaction between the atoms strongly affects entanglement in the system overall and increases the entanglement between the atoms.
+As mentioned above, in reference [28], entanglements between the atoms and the field in the TC model were studied in terms of tangles instead of the von Neumann entropy and entanglement of formation shown in Figures 8 and 10.
+In Figure 11 in D we show tangles corresponding to the system in Figure 10 for $\bar{n}=50$ and $\bar{n}=100$.
+The curves for $\bar{n}=50$ in Figure 11A closely resemble the corresponding curves in Figure 10F, showing that our conclusions are independent of the particular entanglement measure.
+The curves for $\bar{n}=100$ in Figure 11B exactly reproduce the corresponding curves in Fig.
+2a in [28].
 Our contribution aims to discuss the basic phenomenon of two Rydberg atoms interacting with a quantum single-mode electromagnetic field under perfect Rydberg blockade conditions, allowing for an essentially analytical treatment.
@@
 Furthermore, recent successful experimental realizations of photon-photon quantum gates use electromagnetically induced transparency to map Rydberg interactions to light [69, 42]; a realistic description of light-matter interactions of this type in the quantum regime will eventually merge quantum optics with a many-body quantum theory of condensed matter physics, where the Jaynes-Cummings model as described here may serve as a reference point [40, 13].
-6 Acknowledgements
+Figure 10: Time-dependence, in units of $\lambda^{-1}$, of the atomic probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)$, $P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)$, and $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ (see (126) in D), as well as the three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ defined in (70), (73), and (78), respectively, obtained for the TC model in (79) (compare Figure 8).
+The sum of the atomic probabilities is equal to 1 by normalization.
+The parameters, labels and the meaning of the colored lines are the same as in Figure 8; the only exception are the probabilities $P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)$ shown as purple lines in parts (A), (C), (E), which are absent in Figure 8 because the doubly excited state |$\ket{rr}$ is not accessible for Rydberg-interacting atoms.
+As in Figure 8, results are shown for $\bar{n}=10$, $20$, and $50$ for $\omega_{f}=\omega_{0}=\lambda$, zero detuning ($\Delta=0$), and for the initial state |$\ket{g,g,\alpha}$.
+See text and D.
+Acknowledgements
 We thank Volker Quetschke for helpful comments and a critical reading of the manuscript. This work was supported by NSF grants OMA-2231387 and PHY-2012172.
 Appendix A Derivations for Section 3.2: Case B
-In this appendix we first derive results for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ discussed in Section 3.2 (Appendix A.1), and then for the complementary initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2).
+We first derive results for the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$ discussed in Section 3.2 (A.1), and then for the complementary initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (A.2).
 A.1 Initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}=\ket{g,g,n+1}$
@@
 $=\frac{1}{\sqrt{2}}\sin(\varphi_{n})\cos(\varphi_{n})\left[\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)-\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)\right]$ (85)
-$=\frac{1}{\sqrt{2}}\,\frac{1}{2}\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\exp\left(-i\Omega_{n}t\right)-\exp\left(i\Omega_{n}t\right)\right]$ (76)
-$=\frac{1}{\sqrt{2}}(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,,$ (77)
+$=\frac{1}{\sqrt{2}}\,\frac{1}{2}\sin(2\varphi_{n})\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\exp\left(-i\Omega_{n}t\right)-\exp\left(i\Omega_{n}t\right)\right]$ (86)
+[…]
 $\nu_{B}^{(n)}(t)$ $=\left\langle\psi_{2}^{(n)}\Big{|}\psi_{B}^{(n)}(t)\right\rangle=\mu_{B}^{(n)}(t)\,,$ (88)
@@
 $=\sin^{2}(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{+}^{(n)}t\right)+\cos^{2}(\varphi_{n})\exp\left(-\frac{i}{\hbar}E_{-}^{(n)}t\right)$ (91)
-$=\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\sin^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\cos^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,.$ (82)
+$=\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\sin^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\cos^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,.$ (92)
 Equations (87) and (88) imply for the corresponding probabilities
@@
 A.2 Initial state |$\ket{\psi_{\beta}^{(n)}(0)}=\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$
-For later use (Appendix B), here we quote results for the coefficients $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ according to (36) and the resulting probabilities for the initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}$, which by (46) and (87) is complementary to the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}$ in Appendix A.1.
-Similarly as in Appendix A.1 one finds
+For later use (B), here we quote results for the coefficients $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ according to (37) and the resulting probabilities for the initial state |$\ket{\psi_{\beta}^{(n)}(0)}:=\ket{\text{sym},n}$, which by (47) and (97) is complementary to the initial state |$\ket{\psi_{B}^{(n)}(0)}=\ket{\psi_{3}^{(n)}}$ in A.1.
+Similarly as in A.1 one finds
 |$\ket{\psi_{\beta}^{(n)}(t)}$ $=\exp\left(-\frac{i}{\hbar}\hat{H}t\right)\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (98)
@@
 Combining (100) with (20), (21) one obtains
-$\mu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (91)
-$=\frac{1}{\sqrt{2}}\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\left[\cos^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\sin^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,,$ (92)
+$\mu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{1}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=\frac{1}{\sqrt{2}}\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\left[\cos^{2}(\varphi_{n})\exp\left(-i\Omega_{n}t\right)+\sin^{2}(\varphi_{n})\exp\left(i\Omega_{n}t\right)\right]\,,$ (101)
 $\nu_{\beta}^{(n)}(t)$ $=\left\langle\psi_{2}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=\mu_{\beta}^{(n)}(t)\,,$ (102)
-$\xi_{\beta}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle$ (94)
-$=(-i)\sin(2\varphi_{n})\exp\Big{[}-i\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}t\Big{]}\sin(\Omega_{n}t)\,.$ (95)
+$\xi_{\beta}^{(n)}(t)$ $=\left\langle\psi_{3}^{(n)}\Big{|}\psi_{\beta}^{(n)}(t)\right\rangle=(-i)\sin(2\varphi_{n})\exp\Big{[}-i\hskip 1.0pt\Big{(}\omega_{f}n+\tfrac{\Delta}{2}\Big{)}\hskip 1.0ptt\Big{]}\sin(\Omega_{n}t)\,.$ (103)
 The probability $P_{\beta}^{(n)}(\text{sym},t)$ of finding the system in the state |$\ket{\text{sym},n}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ results as, using |$\ket{\psi_{\beta}^{(n)}(t)}=\sqrt{2}\,\mu_{\beta}^{(n)}(t)\ket{\text{sym},n}+\xi_{\beta}^{(n)}(t)\ket{\psi_{3}^{(n)}}$,
@@
 Appendix B Derivations for Section 3.3: Case C
-For the initial states |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ considered in Section 3.3 and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ the time evolution |$\ket{\psi^{(n)}(t)}$ can be deduced from case A (Section 3.1) and Appendix A.2, using
+For the initial states |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}=\ket{r,g,n}$ considered in Section 3.3 and |$\ket{\psi_{2}^{(n)}}=\ket{g,r,n}$ the time evolution |$\ket{\psi^{(n)}(t)}$ can be deduced from case A (Section 3.1) and A.2, using
 |$\ket{\psi_{1}^{(n)}}$ $=\frac{1}{\sqrt{2}}\left(\ket{\psi_{\beta}^{(n)}(0)}+\ket{\psi_{A}^{(n)}(0)}\right)$ (106)
 |$\ket{\psi_{2}^{(n)}}$ $=\frac{1}{\sqrt{2}}\left(\ket{\psi_{\beta}^{(n)}(0)}-\ket{\psi_{A}^{(n)}(0)}\right)\,,$ (107)
-with |$\ket{\psi_{\beta}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (Appendix A.2) and |$\ket{\psi_{A}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (Section 3.1).
+with |$\ket{\psi_{\beta}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (A.2) and |$\ket{\psi_{A}^{(n)}(0)}=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (Section 3.1).
 For the initial state |$\ket{\psi_{C}^{(n)}(0)}=\ket{\psi_{1}^{(n)}}$, (106) implies for the coefficients in (37)
@@
 Using $\mu_{\beta}^{(n)}(t)$, $\nu_{\beta}^{(n)}(t)$, $\xi_{\beta}^{(n)}(t)$ in (101) - (103) and $\mu_{A}^{(n)}(t)$, $\nu_{A}^{(n)}(t)$ in (40) results in (51) - (53) quoted in Section 3.3.
-Appendix C Derivation of $\gamma(t)$ and $\left|\gamma(t)\right|^{2}$ in (68), (69)
-$\gamma(t)$ $=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\langle\text{sym}\,|\,\text{tr}_{f}\left\{\hat{\rho}(t)\right\}|\,g,g\rangle$ (103)
+Appendix C Derivation of $\gamma(t)$ in (72)
+$\gamma(t)$ $=\langle\text{sym}\,|\,\hat{\rho}_{\text{atoms}}(t)\,|\,g,g\rangle=\langle\text{sym}\,|\,\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}(t)\right\}|\,g,g\rangle$ (111)
 $=\sum\limits_{k=0}^{\infty}\left\langle\text{sym},k\,\Big{|}\psi^{(\alpha)}(t)\right\rangle\left\langle\psi^{(\alpha)}(t)\Big{|}\,g,g,k\right\rangle$ (112)
@@
 The result (115) is obtained by using the expressions for $\Delta=0$ of the coefficients $\sigma_{B}^{(n)}(t)=\sqrt{2}\mu_{B}^{(n)}(t)$ in (43) and $\xi_{B}^{(n)}(t)$ in (46).
-Using (67) and the normalization condition $P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)+P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)=1$ one obtains
-$\rho_{\text{atoms}}^{2}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)^{2}+\left|\gamma(t)\right|^{2}&\gamma(t)\\[4.0pt] \gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{g,g},t\right)^{2}+\left|\gamma(t)\right|^{2}\end{pmatrix}$ (109)
-which yields (69). For $\left|\gamma(t)\right|^{2}$ we only quote the result for $\Delta=0$ (as used in Figure 7) for simplicity, using (107):
-$\left|\gamma(t)\right|^{2}=\exp\left(-2\bar{n}\right)\bar{n}\left[\,\sum\limits_{k=0}^{\infty}\frac{\bar{n}^{k}}{k!}\frac{1}{\sqrt{k+1}}\sin\left(\lambda\sqrt{2(k+1}\,t\right)\cos\left(\lambda\sqrt{2k}\,t\right)\right]^{2}\,,\,\,\Delta=0\,,$ (110)
-where $\bar{n}=\left|\alpha\right|^{2}$.
+Appendix D Tavis-Cummings model
+Here we briefly review results for the two-atom Tavis-Cummings (TC) model described by the Hamiltonian (79) as needed to derive the probabilities and entanglements for the TC model shown in Figure 10, according to the notation and definitions used in Sections 2 - 5.
+These results can be deduced from previous work, see, e.g., references [28, 81].
+The total Hilbert space $\mathbb{H}$ of the TC model (79) decouples into 4-dimensional subspaces
+$\mathbb{H}_{n}=\text{span}\left\{\ket{r,r,n-1},\ket{r,g,n},\ket{g,r,n},\ket{g,g,n+1}\right\}\,,\,\,n=1,2,\ldots$ (117)
+including the doubly excited atomic state |$\ket{r,r,n-1}$ (compare the text in the second paragraph below equation (1) in [28]).
+As in Section 2, the index $n$ for $\mathbb{H}_{n}$ corresponds to the number of photons in the states |$\ket{r,g,n}$ and |$\ket{g,r,n}$ (see text below (11)); for $n=0$, the Hilbert space is spanned by the 3 states |$\ket{r,g,0}$, |$\ket{g,r,0}$, |$\ket{g,g,1}$, and for the index $n=-1$ only the state |$\ket{g,g,0}$ is available to the system.
+In the basis corresponding to (117) the Hamiltonian ${\hat{H}}_{TC}^{\,(n)}$ is represented by a $4\times 4$ matrix
+${\hat{H}}_{TC}^{\,(n)}=\hbar\begin{pmatrix}n\omega_{f}&\lambda\sqrt{n}&\lambda\sqrt{n}&0\\[4.0pt] \lambda\sqrt{n}&n\omega_{f}&0&\lambda\sqrt{n+1}\\[4.0pt] \lambda\sqrt{n}&0&n\omega_{f}&\lambda\sqrt{n+1}\\[4.0pt] 0&\lambda\sqrt{n+1}&\lambda\sqrt{n+1}&n\omega_{f}\end{pmatrix}\,\,,\,n=1,2,\ldots$ (118)
+For the initial state |$\ket{\psi^{(n)}_{TC}(0)}=\ket{g,g,n+1}$ at $t=0$ (corresponding to case B discussed in Section 3.2) the time evolution of the system is given by
+|$\ket{\psi^{(n)}_{TC}(t)}$ $=\exp\left(-\frac{i}{\hbar}\hat{H}_{TC\,}t\right)\ket{g,g,n+1}$
+$=\rho_{n}(t)\ket{r,r,n-1}+\sigma_{n}(t)\ket{\text{sym},n}+\xi_{n}(t)\ket{g,g,n+1}\,\,,\,n=1,2,\ldots$ (119)
+where
+$\rho_{n}(t)$ $=\exp(-i\omega_{f}nt)\,a(n)\,b(n)\left(\cos\left(\Omega_{n}^{TC}t\right)-1\right)\,,$ (120)
+$\sigma_{n}(t)$ $=-i\exp(-i\omega_{f}nt)\,a(n)\sin\left(\Omega_{n}^{TC}t\right)\,,$ (121)
+$\xi_{n}(t)$ $=\exp(-i\omega_{f}nt)\left(a(n)^{2}\cos\left(\Omega_{n}^{TC}t\right)+b(n)^{2}\right)\,,$ (122)
+with $a(n)=\sqrt{(n+1)/(2n+1)}$, $b(n)=\sqrt{n/(2n+1)}$, and the Rabi frequency of the Tavis-Cummings model  [^10] Note that $\Omega_{n}^{TC}$ in (123) is different from the Rabi frequency $\Omega_{n}(0)$ in (30) by an extra factor of 2 in front of $n$.
+$\Omega_{n}^{TC}=\lambda\sqrt{2(2n+1)}\,\,.$ (123)
+The normalization $\left|\hskip 1.0pt\rho_{n}(t)\right|^{2}+\left|\hskip 1.0pt\sigma_{n}(t)\right|^{2}+\left|\hskip 1.0pt\xi_{n}(t)\right|^{2}=1$ is fulfilled due to $a(n)^{2}+b(n)^{2}=1$.
+If the two atoms are initially in their ground states |$\ket{g}$ and the field is in a quantum single-mode coherent state |$\ket{\alpha}$ as in (56), the initial state |$\ket{\psi_{TC}^{(\alpha)}(0)}=\ket{g,g,\alpha}$ evolves to
+|$\ket{\psi_{TC}^{(\alpha)}(t)}=\exp\left(-\frac{\left|\alpha\right|^{2}}{2}\right)\left[\exp(i\omega_{0}t)\ket{g,g,0}+\sum\limits_{n=0}^{\infty}\frac{\alpha^{n+1}}{\sqrt{(n+1)!}}\ket{\psi^{(n)}_{TC}(t)}\right]\,,\,\,t>0\,,$ (124)
+with |$\ket{\psi^{(n)}_{TC}(t)}$ in (D) (compare (59)).
+The reduced density operator for the atoms, $\hat{\rho}^{TC}_{\text{atoms}}(t)$, is given by the partial trace of the density operator $\hat{\rho}_{TC}(t)$ of the total system (atoms + field) in the pure state |$\ket{\psi_{TC}^{(\alpha)}(t)}$ over the field |$\ket{\phi_{f}}$, i.e.,
+$\hat{\rho}^{TC}_{\text{atoms}}(t)=\text{tr}_{f}\left\{\hskip 1.0pt\hat{\rho}_{TC}(t)\right\}=\sum\limits_{k=0}^{\infty}\langle k\,|\,\psi_{TC}^{(\alpha)}(t)\rangle\langle\psi_{TC}^{(\alpha)}(t)\,|\,k\rangle\,,$ (125)
+where |$\ket{k}$, $k=0,1,2,\ldots$ (compare (62) and (63)).
+In the basis {|$\left\{\ket{rr},\ket{\text{sym}},\ket{gg}\right\}$, with |$\ket{\text{sym}}=\frac{1}{\sqrt{2}}\left(\ket{r,g}+\ket{g,r}\right)$, the reduced density operator for the atoms is represented by the matrix
+$\rho^{TC}_{\text{atoms}}(t)=\begin{pmatrix}P_{\text{atoms}}^{(\alpha)}\left(\ket{rr},t\right)&\theta^{*}(t)&\phi(t)\\[4.0pt] \theta(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{\text{sym}},t\right)&\gamma(t)\\[4.0pt] \phi^{*}(t)&\gamma^{*}(t)&P_{\text{atoms}}^{(\alpha)}\left(\ket{gg},t\right)\hskip 1.0pt\end{pmatrix}\,\,.$ (126)
+[…]
+$\overline{\rho}^{TC}_{\text{atoms}}(t)=\begin{pmatrix}P_{gg}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\frac{1}{\sqrt{2}}\gamma^{*}(t)&\phi^{*}(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{\sqrt{2}}\theta(t)\\[4.0pt] \frac{1}{\sqrt{2}}\gamma(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{2}P_{\text{sym}}(t)&\frac{1}{\sqrt{2}}\theta(t)\\[4.0pt] \phi(t)&\frac{1}{\sqrt{2}}\theta^{*}(t)&\frac{1}{\sqrt{2}}\theta^{*}(t)&P_{rr}(t)\end{pmatrix}\,\,,$ (127)
+where the overbar indicates that the matrix is with respect to the standard basis.
+We use the same type of abbreviation for the probabilities as quoted below (77) with $P_{gg}(t)+P_{\text{sym}}(t)+P_{rr}(t)=1$ by normalization.
+The three types of entanglement $E(ab-f;t)$, $E(a-bf;t)$, and $E(a-b;t)$ shown in Figure 10 for the TC model were calculated using the density matrices $\rho^{TC}_{\text{atoms}}(t)$ and $\overline{\rho}^{TC}_{\text{atoms}}(t)$ in (126) and (127), respectively, as described in Section 5.
+In reference [28], entanglements between the atoms and the field in the TC model were studied in terms of tangles, which is a different entanglement measure than the von Neumann entropy and entanglement of formation discussed in Section 5 and shown in our Figures 8 and 10 (see equations (7), (8), and (2) in [28]).
+In terms of tangles, the entanglement between the two-atom system $ab$ and the field $f$ is given by (instead of $E(ab-f;t)$ in (70))
+$\tau(ab-f;t)=2\left[1-\text{tr}\left\{\left(\overline{\rho}^{TC}_{\text{atoms}}(t)\right)^{2}\right\}\right]\,.$ (128)
+The entanglement between atom $a$ and the (atom $b$)-field system $bf$ is given by (instead of $E(a-bf;t)$ in (73))
+$\tau(a-bf;t)=2\left[1-\text{tr}\left\{\left(\rho^{TC}_{a}(t)\right)^{2}\right\}\right]\,,$ (129)
+where the reduced density matrix $\rho^{TC}_{a}$ of atom $a$ is the partial trace of $\overline{\rho}^{TC}_{\text{atoms}}(t)$ over atom $b$ (compare (74)).
+Finally, the atom-atom entanglement is given by the tangle (instead of $E(a-b;t)$ in (78))
+$\tau(a-b;t)=C\left(\hskip 1.0pt\overline{\rho}^{TC}_{\text{atoms}}(t)\right)^{2}\,,$ (130)
+where $C\left(\hskip 1.0pt\overline{\rho}^{TC}_{\text{atoms}}(t)\right)=\max\left\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\right\}$ is the concurrence for the density matrix $\overline{\rho}^{TC}_{\text{atoms}}(t)$ (see Section 5.3; here, 3 of the 4 $\lambda_{i}$’s are nonzero).
+For reference purposes, Figure 11 shows the tangles in (128) - (130) for $\bar{n}=50$ and $\bar{n}=100$ for the same system as considered in Figure 10.
+The curves for $\bar{n}=50$ in Figure 11A closely resemble the corresponding curves in Figure 10F, and the curves for $\bar{n}=100$ in Figure 11B exactly reproduce the corresponsing curves in Fig.
+2a in [28].
+Figure 11: Tangles in (128) - (130) for (A) $\bar{n}=50$ and (B) $\bar{n}=100$ for the same system as considered in Figure 10.